### v1

# Equality Saturation for Optimizing High-Level Julia IR

Abstract.
Compilers are indispensable for transforming code written in high-level languages into performant machine code, but their general-purpose optimizations sometimes fall short. Domain experts might be aware of certain optimizations that the compiler is unable to apply or that are only valid in a particular domain. We have developed a system that allows domain experts to express rewrite rules to optimize code in the Julia programming language. Our system builds on e-graphs and equality saturation. It can apply optimizations in the presence of control flow and side effects. As Julia uses multiple dispatch, we allow users to constrain rewrite rules by argument types, and propagate type information through the e-graph representation. We propose an ILP formulation for optimal e-graph extraction taking into account dominance properties for code reuse and introduce CFG skeleton relaxation to rewrite calls to pure functions as well as those with side effects. Use cases demonstrate that our system can perform rewrites on high-level, domain-specific code, as well as on lower-level code such as Julia’s broadcasting mechanism. Finally, we analyze the required compilation time.
1. Introduction
Optimizing compilers have become indispensable as programming languages rely on them to unravel high-level abstractions into lean and performant code (Lattner et al., 2020; Hoste and Eeckhout, 2008; Lattner and Adve, 2004).
General-purpose compiler middle-ends typically perform general-purpose optimizations such as common subexpression elimination (Muchnick, 1998). Compilers for domain-specific languages perform domain-specific optimizations such as algebraic optimizations (Lattner et al., 2021; Muchnick, 1998). The front-ends of higher-level language compilers can be customized to generate IR that is already optimized for specific targets, such as GPUs, and then fed to the back-ends (Besard et al., 2019). Many compilers can also be configured to enable context-specific optimizations, e.g., to sacrifice precision for performance.
However, none of those existing optimization pipelines is well suited to support domain-specific and context-specific optimization of code written in a high-level scientific programming language that is used across many domains, such as Julia (Bezanson et al., 2012). In such a language, software libraries build on the generic language infrastructure to provide high-level domain-specific APIs to application developers.
Ideally, those application developers should be able to enjoy domain-specific optimizations for free, i.e., without having to manually tune or rewrite their code, and without facing restrictions on their freedom to exploit the language’s rapid prototyping features. Furthermore, it should be easy for application developers to specify or select which context-specific optimization they want to see enabled. In addition, when multiple libraries are being reused and composed in an application, be it top-level libraries that provide orthogonal functionality or higher-abstraction-level libraries that provide more abstract APIs on top of lower-level code, the optimization opportunities coming with those libraries should be composable. They should be composable with each other and with the general-purpose and target-specific optimizations that are provided in the main compiler flow and in the target-specific extensions thereof.
This inevitably involves solving a phase-ordering problem because some optimizations create opportunities for further optimizations, while others are incompatible and rule each other out. Finally, to obtain a thriving ecosystem of domain libraries, the development of such libraries and the corresponding support for domain-specific optimization should not require the involvement of compiler experts.
It then follows that a generic compiler infrastructure is needed that enables developers of domain-specific libraries to specify domain-specific and context-specific optimization opportunities at an abstraction level similar to that of their libraries’ APIs, and that ensures sufficient composability.
With our research, we aim to provide that generic compiler infrastructure. In this paper, we present a system that allows programmers, not necessarily compiler developers, to express their domain-specific knowledge as rewrite rules in Julia. Our work is based in part on ideas developed for the Cranelift compiler (Fallin, 2023), which in turn uses e-graphs and equality saturation, first introduced by Nelson (Nelson, 1980) and Tate et al. (Tate et al., 2009), respectively. Equality saturation is a rewrite technique that elegantly deals with the phase ordering problem by representing the potential results of rewrites while still representing the original program as well.
When writing and executing code, users should not take these rewrite rules into account anymore, but instead can focus on writing simple and readable code that maps well on their mental model of the computations. The compiler then automatically evaluates all relevant rewrites and picks the optimal program. By building on e-graphs and equality saturation, the phase ordering problem is no longer an issue. Rewrite rules are applied without removing information from the e-graph, which means that application of a rewrite cannot prevent other rules from firing.
This paper’s contributions are as follows:
the first use of equality saturation in a high-level, dynamically typed programming language optimizer, including for code containing control flow;
CFG skeleton relaxation to enable rewriting on functions with side-effects;
a novel ILP formulation to enable value reuse in (rewritten) expressions;
a novel e-class analysis to track type information for type-constraint rewrite rules;
a demonstration on a number of examples;
an analysis of the required compilation time.
2. Background
E-Graphs
[…]
On the right an e-graph with four e-class nodes is shown for the expression a times two divided by 2: one e-class for each operator and one for each distinct operand values. On the left, an extension of the e-graph is shown in which the equivalent of a times 2 in the form of a shifted to the left by one is included. This graph has two more nodes for the extra shift operation and the extra operand value of 1, but the e-nodes for the times and shift operations form one e-class because of the equivalence.
The e-graph data structure compactly represents a congruence relation on different expression trees (Nelson, 1980; Willsey et al., 2021). It consists of a collection of e-classes that can each contain multiple e-nodes representing expression trees that are equivalent to each other according to some user-defined equivalence relation. Each e-node can have multiple children, represented by e-classes. As an example, the left side of Figure 1 shows the e-graph representation for the expression . In case represents an integer, a potential equivalent expression for is . We can encode this equivalence by adding a new e-node to the appropriate e-class, as shown on the right side of Figure 1. Because the constant expression 1 is not equivalent to any of the e-classes already present, it is added to a newly created e-class. The strength of e-graphs lies in the fact that both equivalent representations are encoded in the graph at the same time.
Equality Saturation
Besides in theorem solvers (de Moura and Bjørner, 2007; Nelson, 1980; De Moura and Bjørner, 2008), e-graphs are often used for equality saturation (Willsey et al., 2021; Cheli, 2021b; Tate et al., 2009). Equality saturation is a term rewriting technique that applies rewrites not by overwriting the original term but by adding the rewritten term to an e-class that represents the original term. By building on e-graphs for term rewriting, it is possible to represent a much larger amount of rewritten, equivalent expressions without the exponential explosion that can occur with naive approaches. The advantage of not discarding expressions when they get rewritten is that the ordering in which rewrites are applied cannot prevent certain rewrites from firing.
Extraction
Once no more new rules can be applied or some timeout is reached during equality saturation, one typically wants to determine the optimal expression contained in the e-graph (Willsey et al., 2021; Goharshady et al., 2024). In its simplest form, extraction starts from a root e-class where one e-node needs to be picked to be extracted. Recursively, all the child e-classes of the picked e-node need to be extracted as well. The end result is a directed, connected graph of e-nodes.
In many use cases, especially in program optimization, the extracted graph cannot contain cycles. In such cases, more refined extraction techniques are needed. One such technique, which considers additional constraints from our problem domain, is discussed in Section 3.6.
3. A Rewrite-Based Compiler Middle-End
Our system allows programmers to specify custom rewrite rules. These rules are then applied on an e-graph representing the code of a function. Once the rules are applied and the e-graph is saturated, the compiler will try to extract the optimal code embedded within. Some example rules are
Rules have a left-hand and right-hand side consisting of a Julia code expression, where the ~ prefix is used to indicate unbound variables. The :: notation is used to specify the types of variables for which the rule can be applied. sin, cos, and translate are the names of called functions. A rewrite rule can be applied whenever the pattern of function calls, literals, and variables on its left-hand side matches a code fragment.
As shown, the two sides of a rewrite rule are specified as regular Julia code expressions, with the addition of unbound, optionally typed, variables. This syntax is able to cover all rewrites where a single expression tree is replaced with a new expression tree. This syntax does not allow rules matching multi-patterns (Jia et al., 2019; Yang et al., 2021). These patterns can match multiple expression trees and rewrite them each, potentially reusing variables in multiple output expressions.
3.1. Incorporating Rewrites in the Compiler
On top, the Julia compiler pipeline is shown: Julia code is parsed into an AST, which is lowered to Julia IR, for which Julia Codegen produces LLVM IR, which is further compiled by LLVM Codegen. On the bottom, the figure zooms in on additional stages that are executed on the Julia IR. These are macro expansion and control flow destructuring to yield untyped IR, type inference and constant propagation that produce typed IR, inlining to obtain inlined IR and finally SROA and simplification to yield optimized IR.
As in most compiler middle-ends, this optimization will operate on an intermediate representation (IR) that supports optimizations while still preserving high-level language-specific information (Lattner et al., 2020; Li et al., 2021; swiftlang, 2012). In our case, this is the Julia compiler’s SSA-based IR (Bezanson et al., 2012).
Figure 2 shows the flow of the compiler. When a function needs to be executed, the just-ahead-of-time flow is invoked on it. We will call this the current function. The compiler first lowers the current function’s AST into an SSA-based IR. On this IR, it applies type inference, constant propagation, function inlining, and a small number of other basic optimizations. From this Julia IR, the compiler then generates LLVM IR, to which many more optimizations are applied by the LLVM compiler before it generates assembly code for the target architecture.
There are two places where it makes sense to apply our rewrite optimizations: right before inlining, and after inlining and optimizations. Before inlining, the current function’s IR still contains function calls as they were written by the programmer. At this stage in the compilation, high-level rewrite rules can typically be matched and applied, i.e., rules involving functions of high-level abstract APIs.
In some cases, it might be useful to also apply rewrites to code in which function calls have already been inlined. After inlining, it is possible that code from different function bodies down the call chain of the current function ends up in its IR. This can uncover additional rewrite opportunities.
Our current prototype implementation performs the rewriting optimization both before and after inlining, but we note that the most interesting rewrites fire before inlining, when high-level API usage has not yet vanished from the code. For this reason, the example snippets of IR in this paper contain code that has not been inlined yet.
3.2. From IR to E-Graph and Back Again
Top left: a simple Julia function that computes x to the power n and that prints a message to illustrate the impact of side effects. Top right: the corresponding Julia IR that is generated if the power function is generated for two integer arguments. It consists of four basic blocks, one of which includes two phi functions. Bottom left: the e-graph for the expressions in the function consisting of 11 e-classes, including classes for the phi functions and for the side effect. Bottom right: the CFG skeleton of the function, which looks like the CFG itself, and in which each node refers to certain e-classes in the e-graph where the control flow depends on expressions modeled by those e-classes.
To apply rewrites on the Julia IR of a function, we first convert it to an e-graph. For this, we employ the methods introduced in the Cranelift compiler (Fallin, 2023). Each function call is added to an e-graph, while control flow statements (goto, goto if not, phi, and return) are added to an external data structure called the CFG skeleton. Function calls that have side effects are also added to the CFG skeleton. Each statement in the CFG skeleton refers to the e-nodes on which it depends. Figure 3 shows an example Julia function pow together with its corresponding IR, e-graph, and CFG skeleton. In addition to the control flow statements, the call to println has been stored in the CFG skeleton because calling that function produces side effects. In this representation, the phi nodes are leaf nodes. Their dependencies are not included in the e-graph. This implies (i) that the e-graph does not contain sufficient information to reconstruct all control flow of the original code, and (ii) that rewrites including phi-node dependencies are not supported. It is the CFG skeleton that provides the extra information needed to reconstruct valid IR.
The need to support control flow stems from a number of common code patterns of expressions that are split over multiple basic blocks for the sake of code conciseness, readability, and maintainability. The most important case concerns if-then-else patterns in which different expressions are computed under different conditions, and of which a common subexpression is computed beforehand. Five idioms for which such patterns occur are the following:
- Keyword Arguments (kwargs):
It is quite common that part of a function depends on a kwarg, e.g., to perform extra checks or to modify the behavior. See, e.g., the use of the flipkernel kwarg in NNLib.jl.111https://github.com/FluxML/NNlib.jl/blob/81e6cd1843dbd2baf2b7b2efd00e942f4f040a2d/src/fold.jl#L186-L191.
- LAPACK-style wrappers with uplo/fmt/trans/… chars:
BLAS/LAPACK-like libraries often use characters to indicate the format of an array (transposed or not, upper/lower triangular, etc.). Such checks are then independent of the values calculated beforehand. For example, in the LinearAlgebra.jl222https://github.com/JuliaLang/LinearAlgebra.jl/blob/master/src/bidiag.jl#L1414-L1448. standard library, the costly operation blks = findall(...) is first computed, after which there is a branch based on uplo, followed by a reuse of blks. Another example is CUDA.jl’s CUSPARSE wrappers: Plenty of API calls (which are a prime use case for e-graph based rewriting) occur outside of fmt checks, others occur inside such checks.333https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusparse/generic.jl#L48-L70.
- API Version Checks:
Continuing on the subject of API calls, it is common that certain API uses are conditional on the version of the library. For example, CUDA.jl includes checks such as CUSPARSE.version() v"11.7.2".444https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusparse/generic.jl#L304-L335.
- Low-level Math:
In scalar ‘mathy’ code, this pattern is also common. The absolute cost of operations outside the if is then often not that great, but when it comes to optimizing very high-performance code, it can make a difference. Some examples can be found in the standard float.jl and math.jl libraries.555https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/float.jl#L320-L328666https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/math.jl#L756-L766.
- Imperative Type Checks:
For readability, or to avoid too many methods, imperative type checks are also sometimes used in Julia instead of multiple dispatch. Several examples can be found in the standard libraries.777https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusolver/linalg.jl#L35-L48, https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/abstractarray.jl#L2485-L2497, https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/intfuncs.jl#L592-L606.
All of these patterns involve divergent control flow, which our work can handle without problems.
To convert an e-graph back to IR, the statements in the CFG skeleton are then used as starting points. This ensures that side effects and control flow are kept intact. IR is created by materializing each statement in the CFG skeleton as well as all its dependencies in the e-graph. Using the scoped elaboration algorithm (Fallin, 2023), materialization occurs only for values that have not yet been materialized in statements that dominate the current statement in the CFG. Otherwise, the previously materialized SSA value is reused.
Like MLIR and the Swift Intermediate Language (SIL) (swiftlang, 2012; MLIR, 2019), Cranelift’s SSA IR format uses basic block arguments, which are bound by the branch operations in a block’s predecessors, to represent control-dependent data flow. Julia’s SSA IR, by contrast, uses phi-nodes. Basic block arguments are more convenient for materializing statements in the e-graph because all information regarding out-of-block uses of expressions that need to be materialized within a block are then explicitly represented within that block itself.
For this reason, our code representation differs from existing uses of e-graphs and CFG skeletons in that it contains both phi-nodes and basic block arguments. Take, for example, the two goto #2 statements in the Julia IR in Figure 3. Since the destination block (#2) contains two phi nodes, the goto statements in the CFG skeleton pass the respective values as arguments. The e-graph itself still stores phi node values, but these are now handled as basic block arguments. During conversion from IR to e-graph and back from e-graph to IR, a conversion between phi nodes and basic block arguments is therefore performed.
3.3. Rules and Types
One of the defining features of the Julia programming language is multiple dispatch. This feature allows users to define a new method of a function for a specific combination of argument types. The fact that the semantics of a function call are not only decided by syntax but also by the type of their arguments necessitates a way to represent this information in rewrite rules as well. For example, a user might want to rewrite matrix multiplications into a function call to some external BLAS library. This means that a call to Base.:*, a function in Julia’s base library, should be rewritten, but only if its arguments are matrices.
Users can specify type constraints in the rewrite rules by associating a type with an unbound variable. For example, a rule for rewriting a matrix multiplication can be written as
We designed a novel e-class analysis for tracking types in the e-graph. For each e-class, this analysis maintains the most specific type that is compatible with all the e-nodes in that class. During e-graph construction, this type is simply the same as the types of statements in the source Julia IR. When two e-classes with a different associated type are merged, the resulting e-class is determined by the typejoin function. For rewrite rules that introduce a new function call, such as the rewrite of a matrix multiplication into blas_call, we run Julia’s type inference using the types stored in the e-classes. Similarly, if the type associated with an e-class changes, we rerun type inference for all e-classes containing e-nodes that depend on the changed e-class.
3.4. Constants in the E-Graph
Rewrite rules are usually symbolic in nature. They match an expression and replace that expression with a new one, plugging in any matched variable. Modern equality saturation frameworks such as egg (Willsey et al., 2021) and Metatheory.jl (Cheli, 2021b), which was used in this work, also support dynamic rewrites. When dynamic rewrite rules are applied, the right-hand side is first executed (potentially using additional e-class analysis data) and the original code is rewritten into the result of that execution, rather than into the right-hand side itself.
When literals are part of the e-graph, dynamic rewrites can be used to perform simple constant propagation. In normal Julia code, literals of primitive types as well as composite types with a known size can appear directly in the IR. The values of these literals are embedded in the e-graph. The e-classes containing the corresponding e-nodes of the literals have an inferred type associated with them, but compared to e-classes containing nonliteral e-nodes, the actual value of that type is available as well. We extend rewrite rule type annotations with a special Comptime{T} type that allows to match against the type of the actual values in the e-graph instead of the inferred type stored in the e-class. Using instead of to signify dynamic rewrite rules, a rule such as
will match any addition between two integer literals and insert the result of this addition in the respective e-class.
Dynamic rewrite rules also open opportunities for partial evaluation-like optimization (Muchnick, 1998; Futamura, 1999) beyond the constant propagation capabilities of the Julia compiler. The compiler does not propagate dynamic array arguments as their values could be mutated at any time. With dynamic rewrite rules, a programmer can override this default behavior. To achieve this, we extended the rewrite rule format with rules of the form
This rule introduces an equivalence between the function’s second argument in its e-graph and the constant W. A constant array holding W’s contents is added as an equivalent e-node to that argument’s e-class in the e-graph, to which dynamic rewrite rules can then be applied. An interesting case is, for example, when W is a matrix holding the constant weights of a neural network layer. Computations that only depend on W or other compile-time-known variables can then be partially evaluated.
3.5. Rewrites in the Presence of Side Effects
We depend on the static effect analysis built in the Julia programming language to decide whether a function call has side effects. This analysis determines several program properties for every statement in the IR of a function. The program property effect_free signals whether a statement is free from externally semantically visible side effects. A limitation of Cranelift’s approach using the CFG skeleton is that there is no way to rewrite instructions with side effects. In Cranelift, this is less of a concern as the code that is rewritten is lower level and typically contains a large amount of simple, low-level instructions, many of which are pure. In high-level Julia code, however, the IR we work on generally is much smaller and consists only of a few calls to high-level functions. Because these functions have more complex behavior than simple, low-level instructions, they also often include behavior with side effects. Still, it should be possible for developers to rewrite function calls that have side effects. For example, a user might want to rewrite an expression containing a function that allocates and returns an array.
To enable this, we introduce a technique called CFG skeleton relaxation. We allow users to create rewrite rules that match against functions with side effects. When such a rule matches an expression in the e-graph and the rewrite is applied by adding the equivalence to the e-graph, all function calls that are part of the matched rule are removed from the CFG skeleton.
However, the ability to ignore the effect analysis and overwrite function calls with side effects is not without risks. Users who are unaware of certain side effects might unknowingly remove them and run into hard-to-debug bugs. We mitigate this problem somewhat by requiring an explicit prefix unsafe in front of the rules that are allowed to remove elements from the CFG skeleton.
Although this raises the bar somewhat to specify and use rewrite rules that unintentionally break the intended semantics of a rewritten piece of code, the risk of using rules incorrectly is obviously not mitigated completely. Semantic rule verification is outside the scope of this work. A rewrite rule might be semantically correct in one domain but incorrect in another, such as reduced precision in deep learning, so correctness is up to the rule author. Bad rules can introduce bugs.
3.6. Extraction
Example function where equality saturation reveals opportunities for reuse. Based on some condition, the function returns either the trace of matrix A or its determinant. The invocations to the functions tr and det that compute the trace respectively the determinant, can be rewritten as a sum and a product over the eigenvalues stored in eig, which is the variable in which the eigenvalues are already stored after a call to the eigen function. The Julia IR is shown next to the source code, and the e-graph is chosen, which consists of 9 e-classes totaling 11 e-nodes. The nine optimal e-nodes, one for each e-class, for extraction are shaded gray.
When Julia IR is generated from an optimized e-graph and its corresponding CFG skeleton, each statement in the CFG skeleton is considered a root node for e-graph extraction. This means that, in contrast to most work on e-graph-based optimizers, we essentially need to run multiple extractions to recover our program. These extractions are not independent of each other. Take, for example, the code and corresponding IR and e-graph in Figure 4. The function f computes and returns two different values eig and x derived from matrix A. eig stores the eigen vectors and values, and x stores the trace or the determinant of the matrix, depending on control flow. In this example, two rules from a collection of linear algebra identities have been applied:
These give rise to the extra e-nodes in the e-classes containing det and tr. Intuitively, we can see that an optimal extraction might be one in which the result of the call to eigen is reused to compute the determinant and/or trace instead of explicitly calling those functions. Indeed, that extraction is valid since the call to eigen dominates both the call to det and to tr so its result is available at those call sites.
We formulate the problem of finding an optimal extraction as an ILP problem. He et al. (He et al., 2017) propose an ILP description for e-graph extraction that ensures an acyclic extracted graph, which is important for further elaboration to produce linearized IR. We adapt and extend their description to take into account e-node reuse from dominating extractions.
An example E-graph with 5 e-classes. E-node n has 2 child e-classes. E-class c contains e-node n, and another, unnamed e-node. Both e-nodes in e-class c have the same two e-node parents. In this graph, nodes are additionally tagged with annotations that illustrate for e-node n the sets p(n) and chi(n).
Let the e-graph be represented by a set of e-nodes , and a set of e-classes . Let be the set of e-nodes that are contained in e-class , and let be the set of children e-classes of a particular e-node . We associate with each e-node a set of e-nodes defined as:
That is, the set of e-nodes that have a child e-class containing . Figure 5 shows an e-graph in which the different parts have been annotated with this notation.
We represent the statements contained in the CFG skeleton as the set . All of these statements cause a different expression to be extracted from the e-graph, rooted at each . For each statement the statements that dominate are represented by the set . Each statement depends on a number of e-class arguments represented by the set .
Finally, we associate a cost with every e-node such that our optimization problem becomes:888Changes to the formulation by He et al. (He et al., 2017) are highlighted in blue.
Where, for each e-node and statement from the CFG skeleton, we have introduced a binary variable . These variables indicate whether the e-node is selected for the extraction rooted at statement . Furthermore, following the work by He et al. (He et al., 2017), each is a class cycle, a collection of classes that contain e-nodes that form a cycle by depending on other classes in the cycle. The e-nodes in an e-class that participate in a class cycle are represented by the set . For each class in each class cycle , we introduce the binary variables that are used to enforce acyclicity in the extracted graph. is assigned 1 if none of the e-nodes in e-class that participate in the class cycle are selected. For a more complete overview and proof of acyclic extraction we refer to the work by He et al. (He et al., 2017). To take into account the dominance information from the CFG skeleton, we adapted their formulation in a few ways.
We adapt constraint (1) to ensure that an e-node can only be selected if for each child e-class an e-node is also selected, or if an e-node is already selected in a dominating extraction .
To prevent e-nodes from being selected in dominating extractions when they are not needed there, we introduce constraint (2). This constraint prevents an e-node from being selected if there is no e-node selected at the same extraction that depends on that e-node.
Constraint (3) ensures that at least one e-node in each e-class that is referenced by a root in the CFG skeleton is truly selected for extraction. Compared to the original ILP formulation by He et al. (He et al., 2017), we added an extra term to account for the fact that an e-node might have already been selected in a dominating extraction.
The remaining constraints are completely analogous to the constraints by He et al. (He et al., 2017) except for the addition that each constraint is added separately for each possible extraction root .
We implement this ILP formulation in JuMP (Lubin et al., 2023), a domain-specific language embedded in Julia for expressing and solving mathematical optimization problems. For a solver, we resort to HiGHS (Huangfu and Hall, 2018). By default, our cost function currently defaults to 1, reducing the optimization problem to choosing the least number of e-nodes. We allow users to manually alter the cost for calls to particular functions to overwrite this behavior.
4. Use Case Demonstration
To demonstrate the capabilities of our work, this section presents a number of use cases. This demonstration focuses on our main goal of allowing users and domain experts to express and exploit optimization opportunities in a convenient, flexible way. In other words, our use cases aim to demonstrate that our work can give developers the necessary control over optimizations without them having to become compiler experts and without restricting the flexibility and productivity they have come to expect from the Julia ecosystem. In contrast, our work aims to give developers more flexibility, meaning less restrictions, to write code without incurring a reduction in optimization opportunities.
Our work does not at all aim to squeeze more performance out of the Julia programming language and its compiler and run-time implementation. Indeed, any optimization that can be obtained with a rewrite rule can also be obtained by manually rewriting source code, such as by duplicating source code and then special-casing it. This section will hence not evaluate how much the performance of a piece of code can be improved with rewrite rules.
4.1. Domain Specific Rewrites
Example of domain-specific code used to apply 2D transformations to a point. The function performs a series of transformations: a horizontal translation over distance dx, a rotation over pi radians, a translation over cosinus squared plus sinus squared minus one, which equals zero, another rotation over pi radians, and a translation over minus dx. With a handful of rules, these can be reduced to the identity transformation.
Figure 6 shows a function that applies a sequence of similarity transformations on a 2D point. On close inspection, application of geometric properties, trigonometric identities, and algebraic simplifications can lead to the reduction of all these transformations to a single identity transform. Indeed, the translation on line 8 can be simplified to p by applying
After this translation has been eliminated, the two rotations by radians similarly end up not transforming the point. Lastly, the first and last rotations now cancel each other out as well. After all the corresponding rewrites, the optimized function simply returns its argument p.
With this example, we also illustrate the necessity of CFG skeleton relaxation. The effect inference in the Julia compiler is unable to reason through the rotation matrix construction in the rotate function on line 2 in Figure 6. Because of this, the rewrite rule author is required to allow CFG skeleton relaxation for those rules where rotate is called by prefixing those rules with unsafe.
4.2. Aiding Multiple Dispatch
Instead of using rewrites purely for simplifying code, they can also be used to generate code that provides additional information to the Julia compiler’ type system to trigger the execution of better performing function implementations through multiple dispatch. An example in the domain of linear algebra is to automatically wrap certain matrix expressions in a new, more specific type that allows more efficient implementations for subsequent computations. Take, for example, the expression between two regular matrices and . For the case where and are the same matrix, that is, , we know that the result is a symmetric matrix. We can encode this fact by applying the following rewrite rule:
The Symmetric function from the LinearAlgebra.jl package takes the upper triangle of its argument and uses that to efficiently represent a symmetric matrix. More importantly, it returns a value of type Symmetric. Other functions in that linear algebra package, e.g., for solving eigenproblems, have specialized implementations for symmetric matrices. After rewriting with the above rule, those specialized implementations will now be called automatically thanks to the multiple dispatch system knowing that they are invoked on type Symmetric instead of on a more generic matrix type.
Another example is the expression , which is symmetric if B is symmetric as well. Again, this fact can be encoded with the following rule:
Both these examples build on the abstractions of the Julia standard library, but the same concepts can also be applied to operations and types in other, external packages.
4.3. Multi-Line Broadcast Fusion
Julia code that uses broadcasting syntax: using the dot operator, a elementwise ReLu operation and an elementwise addition are applied to vectors A and B. In other words, the relu function is broadcasted over matrix A and vector B is added to the result using broadcasting addition.
Julia code example of materialize(broadcasted(+,broadcasted(relu,A),B)) to illustrate syntactic lowering applied for the broadcasting expression in the previous figure showing a broadcasting expression. The broadcasted function builds a lazy representation of the broadcast operation. The materialize function allocates the result array and computes its values.
Julia’s broadcasting mechanism allows users to apply functions element-by-element on one or more arguments containing multiple elements. Syntactically, this is done by adding a . between the function and its argument list. The top of Figure 7 shows two broadcast operations. relu.(A) applies the scalar relu function to each element of the matrix A. Next, the .+ call computes the addition between the result and a vector B. Although the two arguments to .+ have a different shape, the operation can take place because the shapes are compatible.
Broadcasting in Julia is implemented by transforming expressions containing this special dot syntax into calls to the standard library functions broadcasted and materialize. Figure 8 shows the result of this transformation on the broadcasting expression in Figure 7. A call to broadcasted builds a lazy representation of the broadcast result. Different lazy broadcast objects can be nested; it is only when materialize is called that the final result object is allocated and filled with computed values. As such, different operations within multiple broadcasted calls are fused when materialize is called. This leads to less memory being used because temporary arrays are not materialized and potentially better run-time performance because there are fewer function calls. This operator fusion is especially beneficial for code that is executed on the GPU, as kernel launches can take a significant amount of time, and because kernel fusion can significantly reduce the required number of memory accesses (Besard et al., 2019).
A limitation of Julia’s broadcasting lowering is that fusion cannot occur across broadcast expressions on different lines. This is because replacing the dot syntax with calls to broadcasted and materialize is a syntactical transformation that is performed during AST lowering, which operates statement by statement. A potential optimization is hence to eliminate superfluous calls to materialize. Rewrite rules make this easy. Take for example the rule
This rule will remove the intermediary materialization for any call of the form f.(x) where x is a variable that was defined on a different line with a broadcast expression itself. In the example code in Figure 9, fusion will now be applied despite the two dot operators occurring in two different statements on two separate lines. This example illustrates how the system we propose not only supports domain-specific optimization, but at the same time enables optimization of common code patterns involving only core Julia primitives.
An important detail is that here, once more, CFG skeleton relaxation is required, since materialize ends up calling a foreign C function function which the compiler cannot deem effect-free, poisoning the remainder of the effect analysis. In practice, this means the rule needs to be prefixed with unsafe.
Using PkgEval.jl, a tool to automatically evaluate the tests of Julia packages, we evaluated how many times the above rule’s pattern occurs in real-life code in 3153 different packages. In total, we found 514 instances where code containing this pattern was executed, spread over 232 different packages. This pattern occurs, for example, when a broadcasted expression is used in different branches and is factored out by the programmer as a code simplification measure. Applying the rewrite rule does not undo the deduplication but only removes the superfluous call to materialize. Note that the discussed rule only matches expressions where the outer broadcasted function is a unary function.
For other patterns, e.g., an expression where the outer broadcasted function is a binary function, separate rules need to be written. Table 1 gives an overview of how many times different multi-line broadcasting patterns were found in our analysis. The three last patterns are different configurations of broadcasting a binary function. Note that the matches for these three patterns are not independent: a match against the last pattern, which will result in two invocations of materialize being optimized out, implies the two other patterns can be matched and applied as well, each optimizing out only one materialize invocation. When the final code is extracted in such cases, the expression from the e-class containing the least materialize calls will be chosen.
| LHS pattern of rewriting rule | Count | Distinct Packages |
| broadcasted(~f, materialize(~x)) | 514 | 232 |
| broadcasted(~binop, materialize(~x), ~y) | 1232 | 348 |
| broadcasted(~binop, ~x, materialize(~y)) | 1044 | 283 |
| broadcasted(~binop, materialize(~x), materialize(~y)) | 231 | 93 |
4.4. Extending Broadcasting with Domain-Specific Batched Operations
Julia code example of a function called forward that computes the forward pass of a single-layer neural network. Apart from network inputs, weights, and activation function, the forward function takes a boolean argument called training, and an argument called p_drop which is the probability value used in dropout operation. The function starts by assigning the logits variable, which is computed by a matrix multiplication between the weights and input, and a broadcasted addition with the bias vector. Using an if statement, depending on whether the training flag is set true or false, the dropout function is applied before broadcasting the activation function on the result and returning it.
Julia code of the same neural network computation as the previous figure, after rewrite transformations. Now, the branch where no dropout is applied directly a specialized function called fused_dense_bias_activation. The branch where dropout is applied still first computes the logits variable, applies the dropout function, and returns the value after applying the activation function.
In high-performance contexts, programmers often sacrifice code simplicity and readability in pursuit of achieving the highest possible performance. For example, instead of writing deep learning layers and non-linear activation functions using core language constructs, programmers often resort to using frameworks or libraries that contain optimized implementations of matrix multiplications, activation functions, attention computation, and others (Elshawi et al., 2021; Parvat et al., 2017). This works well for most common use cases but falls apart when a programmer requires custom behavior that is not yet supported by the framework. In those cases, the programmer is required to rewrite parts of the program to stop using the framework and instead depend on a different framework or the core language.
Instead of having to opt out of framework functionality and rewriting code, it would be much better if one could write simple, readable code and have the frameworks themselves automatically discover opportunities to use their better performing implementations. Our rewrite system gives framework authors the option for implementing such discovery and optimization almost for free.
As an example, take the code in Figure 9 that shows an implementation of the forward pass of a single fully connected feedforward layer. There already exist several Julia packages that allow us to compute such operation more efficiently (Faingnaert et al., 2022; Innes et al., 2018; Innes, 2018). For example, the programmer could rewrite the code as shown in Figure10 to use the optimized fused_dense_bias_activation function from the Lux.jl deep learning framework (Pal, 2023). A disadvantage of such manual rewriting is that the code becomes less readable for someone not familiar with the Lux.jl APIs. Moreover, the use of that specific framework’s optimized implementation, which is in essence a decision about a non-functional aspect of the code, is then hard-coded in the code that specifies the functionality. This requires the developers and the maintainers of this code to be knowledgable in the code’s application domain as well as in the Lux.jl framework. Switching to other frameworks in the future will require rewriting code. Clearly, hard coding the dependence on Lux.jl and its APIs in the code has a number of downsides.
With our system, instead of rewriting the original code of Figure 9, it suffices to add a rewrite rule that maps the original code pattern onto the optimized fused_dense_bias_activation function:999Type annotations have been left out for the sake of conciseness of this rule, but they are needed to make sure to only match when W and x are matrices, and b is a vector.
| materialize(broadcasted(~, broadcasted(+, ~W * ~x, ~b))) | ||
Combined with the multi-line broadcast rewriting rule explained above, which rewrites the original code into a single line without the intermediary variable logits, this rule maps the single-line broadcast expression onto a call to the library function LuxLib.fused_dense_bias_activation, which will yield compiled code equivalent to what would be obtained with the code in Figure 10.
This example illustrates how our approach allows for a better separation of concerns. Another advantage is that alternative rewrite rules can be provided for different Julia packages that provide different optimized implementations for similar functionality. Different rewrite rules also don’t obfuscate each other. All rewrite rules can be represented in the e-graph. Only during extraction is the final optimized code materialized, optionally based on user-defined extraction costs for different functions. As such, rewrite rules from different authors are fully interoperable.
Similarly, without changing the source code, different rewrites can still occur by applying different sets of rewrite rules that can for example be offered by different Julia packages.
This example also illustrates how our system supports the composition of high-level, domain-specific optimizations with optimizations of core Julia primitives. In this way, our system’s scope is much broader than recent equality saturation based optimizers such as Cranelift.
5. Compilation Time Evaluation
Julia is a just-ahead-of-time compiled language for scientific computing. By default, Julia automatically compiles a native, type-specialized version of a function just ahead of the first time it is to be executed on some combination of argument types. Julia hence positions itself somewhere between typical ahead-of-time compilers such as LLVM and gcc and just-in-time compilers such as Cranelift or the compilers used in a Java VM. Compilation speed is hence important for Julia.
Alternatively, a user can request ahead-of-time compilation of functions and optionally invoke our rewriting rule-based optimization, depending on whether or not the investment in compilation time might yield a positive return on investment. As scientific software most often has relatively long execution times, compilation times are therefore not as critical as in more traditional JIT compilers. Still, we aim for interactive compilation speeds with our work to allow for rapid prototyping.
5.1. Complexity of E-Graph Construction
In general, termination for equality saturation is not guaranteed since application of rewrite rules can introduce new rewrite opportunities indefinitely. (Willsey et al., 2021; Suciu et al., 2025) In practice, this issue is most often tackled by equality saturation timeouts in combination with heuristic schedulers that ensure that all rewrite rules get a chance to be applied by implementing a rule application back-off mechanism (Cheli, 2021a; Yang et al., 2021). The non-terminating behavior of equality saturation is due to repeated application of rule application and rebuilding the e-graph by joining newly discovered equivalences. The actual construction of the e-graph, however, is not concerned with this loop and does, in fact, execute in polynomial time.
We reimplemented the algorithm introduced for Cranelift (Fallin, 2023) in Julia. This algorithm starts by computing the dominator tree of the input SSA code. We reuse the dominator tree construction implementation from the Julia compiler. This implementation is based on an algorithm from Georgiadis’s thesis (Georgiadis, 2005), and runs in linear time with respect to code size. IR statements are visited in a pre-order dominator tree traversal to ensure that all dependencies of the current statement have already been visited. For each statement, the e-class dependencies are looked up and an e-node is constructed. If the e-node is not part of the e-graph yet, it is inserted in a new e-class, otherwise an existing e-class is associated with the statement. At the same time that the e-graph is being constructed, the CFG skeleton is being built. This involves pushing the corresponding e-classes of effectful statements in a CFG structure which has a worst case runtime complexity of , where is the code size, but an amortized complexity of .
Since no equalities are introduced yet during e-graph construction time, inserting new e-classes in the e-graph does not require running the relatively expensive rebuilding procedure. As such, e-graph construction only requires inexpensive hash map lookups and insertions for each statement, which have a worst case runtime complexity of but an expected worst case runtime of .
Overall, the worst case runtime complexity of our e-graph construction is but on average and in practice, as will be discussed in Section 5.2, the worst case runtime complexity is .
5.2. Experimental Compilation Time Evaluation
A graph showing the compilation time for different parts of our rewrite optimizations. The graph consists of three vertical bar charts next to each other, representing time spent in the pre-inlining, post-inlining, and finalizations stage. Each bar chart is divided in different colors indicating the time needed for specific compilation steps. The different compilation steps are: IR to e-graph conversion, e-graph saturation, e-graph to ir conversion, type reinference, invocation queueing, ILP problem construction, ILP solver (HiGHS), optimized function insertion, inlining pass, llvm optimization and code generation, and other.
Figure 11 shows the compilation times spent in different steps of the whole optimization pipeline for optimizing the forward function from Figure 9. Important to note is that these numbers include the time spent optimizing each function in the call graph of the top-level function. For the forward method, the call graph consists of 43 methods that each pass through the full optimization pass. The chart partitions the optimization flow and its compilation time into three stages. All benchmarks were run on a machine with a 6-core Intel i7-9750H processor and 24GB RAM. The code was run on Julia v1.11.0 (commit 501a4f25c2b), with LLVM 16.0.6.
5.2.1. Pre-Inlining
In this first stage, rewrites are applied before the code body of the function is inlined. Due to the small number of rules, e-graph saturation only takes a small portion of the work in this stage. Similarly, ILP problem construction and solving runs quickly because non-inlined code consists of few statements, leading to a small ILP problem to solve. In this stage, a larger part of the run time is spent reinferring the types in the generated IR. This also includes replacing dynamic function calls with static invocations of the correct methods if enough type information is available to make this decision. The last step of this stage consists of collecting the non-inlined IR of each method that is invoked, and adding this IR to the top of a worklist. Methods invoked within a function end up at the top of the worklist and are fully optimized before the second stage of the current function can begin.
5.2.2. Post-Inlining
The second stage starts by replacing each method invocation with the optimized variant if rewrites were applied to it. Next, we have the default inlining pass that is also invoked during regular Julia compilation. In this second stage, conversion from IR to e-graph takes much longer because the inlined code consists of many more nodes. For the same reason, e-graph saturation takes longer as well. Because there are no rules to apply post-inlining in this instance, no time is spent on ILP, and type re-inference runs faster.
To assess the scalability of the conversion from IR to e-graph and back, Figure 12 shows those times for IRs of different sizes. The IR dataset was collected by running the optimizer on the factorize function from Julia’s linear algebra standard library and on the many functions that are part of this function’s call graph. Both conversions clearly scale linearly with the number of statements.
E-graph to IR conversion does not meaningfully show up in the compilation times in Figure 11 because this operation is only executed if actual changes to the e-graph were made, which happens infrequently in this instance. Similarly, re-inference times in the post-inlining stage do not take much time because this step is only applied when the e-graph has changed. This is in contrast to the pre-inlining stage, where part of the re-inference procedure needs to be executed either way to prepare the IR for inlining in the next stage.
A scatter plot showing the time needed to convert from IR to e-graph and vice versa. The x-axis shows the number of IR statements, the y-axis shows the time needed in microseconds. Datapoints for IR to e-graph conversion are shown with an orange, plus-shaped marker, and e-graph to IR is shown using an red, triangle-shaped marker. Both cases a similar linear trend upwards.
5.2.3. Finalization
During the last stage, the functions that were enqueued in the previous stage have been optimized and are inserted if rewrites were applied to them. Finally, compilation is passed on to LLVM to generate an executable, native function from the optimized IR. This step takes more than a third of the total time.
In Figure 11, Invocation Queueing, Inlining, and LLVM Compilation are steps that are not specific to our compilation flow but are also taken in the native Julia compiler. In the example discussed above, 80% of the total time is spent in these steps. The Julia compiler hides much of the compilation costs by caching the compiled code. Tighter integration of our optimization into the Julia compiler could thus lead to faster compile times as well. As briefly noted in Section 3.1, rewrite rules will be applied more often on IR before inlining because that IR contains function calls roughly as they were written by a programmer. In IR to which inlining has been applied, on the other hand, it is difficult to anticipate which function calls will end up visible, making it less productive to write rules targeting IR at this level. For this reason, a worthwhile trade-off might be to disable rewrites in the second, post-inling stage. This would eliminate the high cost of the IR to e-graph conversion that is included in the middle bar in Figure 11.
5.2.4. ILP Construction and Solving
Table 2 shows the time required to construct and solve the ILP problem for several examples discussed in this paper. With the exception of the code for simplifying 2D transformations in Figure 6, e-classes do not contain many equivalences, leading to fast problem construction and solving. For the 2D transformation example, particular rewrite rules, e.g., those that combine two subsequent rotations into one rotation, give rise to an infinite number of rewrite opportunities. In these instances, saturation runs until a timeout is reached. If fast compilation times are required for pathological cases like the one shown, we can fall back to a greedy extraction algorithm. In the future, better scheduling of rules and detection of these cases could limit the amount of e-graph blow-up as well.
6. Limitations
Control Flow and Inlining
As discussed in Section 3.2, our approach to converting IR to an e-graph using the CFG skeleton currently does not allow applying rewrites that alter control flow. Similarly, rewrite opportunities where part of an expression is hidden behind a function call are in general not considered. We optimize each function separately, applying rewrite rules both before and after inlining is applied. Other intermediate representations, discussed further in Section 7, could be used to represent the entire program in an e-graph and allow rewriting across call boundaries and control flow.
Saturation and Extraction
In this work, we have proposed an ILP formulation for optimal e-graph extraction. The problem of extracting e-nodes from a general e-graph is NP-hard, which means that there is no guarantee that an optimal solution can be found (Goharshady et al., 2024). Currently, when the ILP solver is unable to find a solution, we fall back to a greedy extraction algorithm that is built into Metatheory.jl. In this work, we chose a simple cost model that treats the cost of nodes independent of other nodes. A more complex cost model could take into account more complex behavior that depends on multiple nodes at once.
Rule Expressiveness
As discussed in Section 3, our rule syntax takes one Julia expression and transforms it into another. This is an intuitive syntax for expressing rewrite rules, but does not fully cover all the possible rules a developer could possibly want to express. For one, the rewrite rules can only be used to rewrite a single expression into another expression. Multi-pattern rewrite rules to match against and to rewrite multiple expressions at once are currently not supported. We also do not yet support rewrite rules matching against functions calls taking an unknown, variable amount of variables. This means that in certain cases, a developer has to rewrite multiple similar rules to match the same function with different numbers of arguments.
7. Related Work
7.1. Equality Saturation
Equality saturation and its use as a code optimization tool was first introduced by Tate et al. (Tate et al., 2009), where it was used to optimize Java bytecode. Willsey et al. (Willsey et al., 2021) introduced algorithmic improvements for equality saturation providing asymptotic speed-ups. This has led to an explosion of new work exploring different applications for equality saturation, from optimizing tensor programs (Wang et al., 2020; Yang et al., 2021; Smith et al., 2021), to compiler verification (Stepp et al., 2011; Kourta et al., 2022), optimizations for specialized hardware (Ustun et al., 2022; VanHattum et al., 2021; Matsumura et al., 2023), and other purposes (Panchekha et al., 2015; Chandrakana Nandi et al., 2021).
Metatheory.jl (Cheli, 2021b) provides a framework for equality saturation in Julia. It has already been used for code optimization on symbolic representations of Julia code to speed up PDE solvers (Gowda et al., 2022). That work, however, only handles pure, symbolic programs, as it operates on a symbolic representation of Julia code obtained through manual expression building or tracing through heavily restricted, straight-line code. In contrast to our work, that existing use of Metatheory.jl does hence not allow the presence of general control flow or side effects.
To the best of our knowledge, by introducing type propagation in the e-graph, we are the first to apply equality saturation to a dynamic programming language. This allows rewrites to coexist and complement Julia’s multiple dispatch feature, as discussed in Section 4.
7.2. Intermediate Representations and E-Graphs
Cranelift, a compiler back-end for WebAssembly and Rust, is among the first to use e-graphs and equality saturation in a production-grade compiler (Fallin, 2023). While we reuse ideas and methods first introduced in the Cranelift compiler, we implemented our framework as a standalone project, not reusing any Cranelift components. Our code is integrated in the Julia compiler using user-land compiler extensions.
The Cranelift compiler does not do full equality saturation, but only recognizes equalities the first time a node is inserted in the e-graph. This implies that it is possible for equalities to remain undiscovered, but it obviates the fixed-point loop present in full equality saturation and ensures there are no cycles in the e-graph which allows for a more efficient representation of the graph in memory and leads to easier extraction. We use the same e-graph representation of code as used in Cranelift, but by utilizing an extraction scheme that enforces acyclic extraction, the optimizer is able to carry out full equality saturation. Whereas Cranelift and our work operate on IR that is complemented by a CFG, other work has sought to find alternative IR formats that allow representing control flow fully in the e-graph. Tate et al. (Tate et al., 2009) have introduced Program Expression Graphs (PEGs) which represent constructs such as loops as special expressions and show that PEGs can be used to do equality saturation. Similarly, Regionalized Value State Dependency Graphs (RVSDGs) (Bahmann et al., 2015; Reissmann et al., 2020) represent control flow as (nested) expressions. A research prototype already exists using RVSDGs in conjunction with equality saturation for a simple toy language (Sharp, 2022). Our approach remains closer to the original SSA-based IR of the source language, allowing less expensive conversion routines.
Willsey et al. introduced e-class analyses to associate additional metadata with each e-class (Willsey et al., 2021). The metadata is kept up to date throughout the equality saturation process by potential modifications whenever new information is available, for example, when e-classes are merged. An exemplary existing use case for an e-class analysis is constant propagation, where each e-class can optionally carry a constant value. To track types in the e-graph, we designed a novel e-class analysis, as discussed in Section 3.3
7.3. Extraction
We based our ILP optimal extraction formulation on the work of He et al. (He et al., 2017). Recent work proposes other formulations, for example by looking at e-graphs through the lens of circuits (Sun et al., 2024), or finite state automata (Y. Wang et al., 2022). These formulations can lead to faster extraction or provide termination guarantees under particular assumptions. Cranelift operates in a Just-In-Time context where fast and reliable extraction is crucial. For this reason, they use an extraction algorithm that greedily tries to minimize the total number of nodes, without taking into account node reuse in extracted expressions or from expressions that dominate them. In contrast to Cranelift’s greedy algorithm, we have introduced an ILP formulation that aims for node reuse from other nodes within an extracted expression, as well as from nodes in expressions that dominate them in the CFG.
7.4. Julia Code Optimization
Other works exist that aim to improve the performance of Julia code with custom optimizations. Symbolics.jl (Gowda et al., 2022) can be used to transform scientific machine learning code into a symbolic representation at which point simplifications can be applied. Finch.jl (Ahrens et al., 2024, 2023) uses a custom IR to represent iteration patterns of loop nests over sparse or structured arrays. We instead focus on optimizing code in Julia’s own IR format, allowing users of our system to target any Julia code.
8. Conclusion and Future Work
We have developed a novel system that allows Julia developers to write domain-specific rewrite rules that are automatically applied using equality saturation and that works in the presence of control flow and side effects. Through an e-class analysis, we keep track of the most specific type of all equivalent terms and use this information to support type-constrained rewrite rules. Unlike previous work, our system can rewrite dynamically-typed code. We introduce a technique called CFG skeleton relaxation that allows rewrite rules to nullify side effects in the original code. We have adapted an ILP formulation for optimal, acyclic e-graph extraction to take into account value reuse from dominating statements. Finally, we show how this system can perform rewrites on high-level domain-specific code and how the use of equality saturation solves the phase ordering problem.
In the future, our work can be extended to support rewriting control flow and apply rewrites that span different call depths. Further improvements to our extraction scheme could also lead to more efficient code being generated.
All artifacts will be made available upon acceptance of the paper.

### v2

@@
 Compilers are indispensable for transforming code written in high-level languages into performant machine code, but their general-purpose optimizations sometimes fall short.
-Domain experts might be aware of certain optimizations that the compiler is unable to apply or that are only valid in a particular domain.
+Domain experts might be aware of optimizations that the compiler is unable to apply or that are only valid in a particular domain.
 We have developed a system that allows domain experts to express rewrite rules to optimize code in the Julia programming language.
@@
 As Julia uses multiple dispatch, we allow users to constrain rewrite rules by argument types, and propagate type information through the e-graph representation.
-We propose an ILP formulation for optimal e-graph extraction taking into account dominance properties for code reuse and introduce CFG skeleton relaxation to rewrite calls to pure functions as well as those with side effects.
+We propose an ILP formulation for optimal e-graph extraction that exploits opportunities for code reuse and introduce CFG skeleton relaxation to rewrite calls to pure functions as well as those with side effects.
 Use cases demonstrate that our system can perform rewrites on high-level, domain-specific code, as well as on lower-level code such as Julia’s broadcasting mechanism.
-Finally, we analyze the required compilation time.
+We analyze the required compilation time and the performance impact of these rewrites.
 1. Introduction
-Optimizing compilers have become indispensable as programming languages rely on them to unravel high-level abstractions into lean and performant code (Lattner et al., 2020; Hoste and Eeckhout, 2008; Lattner and Adve, 2004).
+Optimizing compilers have become indispensable as programming languages rely on them to unravel high-level abstractions into lean and performant code (Lattner et al., 2021; Hoste and Eeckhout, 2008; Lattner and Adve, 2004).
 General-purpose compiler middle-ends typically perform general-purpose optimizations such as common subexpression elimination (Muchnick, 1998).
@@
 Many compilers can also be configured to enable context-specific optimizations, e.g., to sacrifice precision for performance.
-However, none of those existing optimization pipelines is well suited to support domain-specific and context-specific optimization of code written in a high-level scientific programming language that is used across many domains, such as Julia (Bezanson et al., 2012).
+However, none of those existing optimization pipelines are well suited to support domain-specific and context-specific optimization of code written in a high-level scientific programming language that is used across many domains, such as Julia (Bezanson et al., 2012).
 In such a language, software libraries build on the generic language infrastructure to provide high-level domain-specific APIs to application developers.
 Ideally, those application developers should be able to enjoy domain-specific optimizations for free, i.e., without having to manually tune or rewrite their code, and without facing restrictions on their freedom to exploit the language’s rapid prototyping features.
-Furthermore, it should be easy for application developers to specify or select which context-specific optimization they want to see enabled.
+Furthermore, it should be easy for application developers to specify or select the context-specific optimization they want enabled.
 In addition, when multiple libraries are being reused and composed in an application, be it top-level libraries that provide orthogonal functionality or higher-abstraction-level libraries that provide more abstract APIs on top of lower-level code, the optimization opportunities coming with those libraries should be composable.
@@
 (Tate et al., 2009), respectively.
-Equality saturation is a rewrite technique that elegantly deals with the phase ordering problem by representing the potential results of rewrites while still representing the original program as well.
-When writing and executing code, users should not take these rewrite rules into account anymore, but instead can focus on writing simple and readable code that maps well on their mental model of the computations.
+Equality saturation, or EqSat, is a rewrite technique that elegantly deals with the phase-ordering problem by representing the potential results of rewrites while still representing the original program as well.
+When writing and executing code, users should not take these rewrite rules into account anymore, but instead can focus on writing simple and readable code that maps well onto their mental model of the computations.
 The compiler then automatically evaluates all relevant rewrites and picks the optimal program.
-By building on e-graphs and equality saturation, the phase ordering problem is no longer an issue.
+By building on e-graphs and EqSat, the phase-ordering problem is no longer an issue.
 Rewrite rules are applied without removing information from the e-graph, which means that application of a rewrite cannot prevent other rules from firing.
-This paper’s contributions are as follows:
-the first use of equality saturation in a high-level, dynamically typed programming language optimizer, including for code containing control flow;
-CFG skeleton relaxation to enable rewriting on functions with side-effects;
-a novel ILP formulation to enable value reuse in (rewritten) expressions;
-a novel e-class analysis to track type information for type-constraint rewrite rules;
-a demonstration on a number of examples;
-an analysis of the required compilation time.
+The contributions in this paper are as follows: (i) the first use of EqSat in a high-level, dynamically typed programming language optimizer, including for code containing control flow; (ii) CFG skeleton relaxation to enable rewriting on functions with side effects; (iii) a novel ILP formulation to enable value reuse in (rewritten) expressions; (iv) a novel e-class analysis to track type information for type-constrained rewrite rules; (v) a demonstration on a number of examples; (iv) an analysis of the required compilation time; (vi) an analysis of the potential performance gains achievable.
+This paper is structured as follows.
+Section 2 provides background.
+Section 3 presents our rewriting system.
+Section 4 presents its capabilities on use cases.
+Section 5 analyzes its compilation time, and Section 6 demonstrates obtainable performance optimizations.
+Sections 7 and 8 discuss limitations and related work.
+Finally, Section 9 draws conclusions.
 2. Background
-E-Graphs
 […]
@@
 This graph has two more nodes for the extra shift operation and the extra operand value of 1, but the e-nodes for the times and shift operations form one e-class because of the equivalence.
-The e-graph data structure compactly represents a congruence relation on different expression trees (Nelson, 1980; Willsey et al., 2021).
-It consists of a collection of e-classes that can each contain multiple e-nodes representing expression trees that are equivalent to each other according to some user-defined equivalence relation.
+E-Graphs.
+An e-graph compactly represents a congruence relation on different expression trees (Nelson, 1980; Willsey et al., 2021).
+It consists of a collection of e-classes that can each contain multiple e-nodes representing expression trees that are equivalent to each other according to a user-defined equivalence relation.
 Each e-node can have multiple children, represented by e-classes.
-As an example, the left side of Figure 1 shows the e-graph representation for the expression .
+As an example, the left side of Fig.
+1 shows the e-graph for the expression .
 In case represents an integer, a potential equivalent expression for is .
-We can encode this equivalence by adding a new e-node to the appropriate e-class, as shown on the right side of Figure 1.
+We can encode this equivalence by adding a new e-node to the appropriate e-class, as shown on the right side of Fig.
+1.
 Because the constant expression 1 is not equivalent to any of the e-classes already present, it is added to a newly created e-class.
@@
 Equality Saturation
-Besides in theorem solvers (de Moura and Bjørner, 2007; Nelson, 1980; De Moura and Bjørner, 2008), e-graphs are often used for equality saturation (Willsey et al., 2021; Cheli, 2021b; Tate et al., 2009).
-Equality saturation is a term rewriting technique that applies rewrites not by overwriting the original term but by adding the rewritten term to an e-class that represents the original term.
-By building on e-graphs for term rewriting, it is possible to represent a much larger amount of rewritten, equivalent expressions without the exponential explosion that can occur with naive approaches.
+Besides in theorem solvers (de Moura and Bjørner, 2007; Nelson, 1980; De Moura and Bjørner, 2008), e-graphs are often used for EqSat (Willsey et al., 2021; Cheli, 2021b; Tate et al., 2009).
+EqSat is a term rewriting technique that applies rewrites not by overwriting the original term but by adding the rewritten term to an e-class that represents the original term.
+By building on e-graphs for term rewriting, it is possible to represent a much larger number of rewritten, equivalent expressions without the exponential explosion that can occur with naive approaches.
 The advantage of not discarding expressions when they get rewritten is that the ordering in which rewrites are applied cannot prevent certain rewrites from firing.
+Central to EqSat is the rebuilding procedure, which is responsible for maintaining the e-graph invariants.
+When rewriting uncovers two different expressions to be equivalent, all expressions containing these equivalent expressions as subexpressions need to be verified, as these potentially also need to become equivalent now.
 Extraction
-Once no more new rules can be applied or some timeout is reached during equality saturation, one typically wants to determine the optimal expression contained in the e-graph (Willsey et al., 2021; Goharshady et al., 2024).
+Once no more new rules can be applied or some timeout is reached during EqSat, one typically wants to determine the optimal expression contained in the e-graph (Willsey et al., 2021; Goharshady et al., 2024).
 In its simplest form, extraction starts from a root e-class where one e-node needs to be picked to be extracted.
@@
 The end result is a directed, connected graph of e-nodes.
-In many use cases, especially in program optimization, the extracted graph cannot contain cycles.
-In such cases, more refined extraction techniques are needed.
-One such technique, which considers additional constraints from our problem domain, is discussed in Section 3.6.
-3. A Rewrite-Based Compiler Middle-End
-Our system allows programmers to specify custom rewrite rules.
-These rules are then applied on an e-graph representing the code of a function.
-Once the rules are applied and the e-graph is saturated, the compiler will try to extract the optimal code embedded within.
-Some example rules are
-Rules have a left-hand and right-hand side consisting of a Julia code expression, where the ~ prefix is used to indicate unbound variables.
-The :: notation is used to specify the types of variables for which the rule can be applied.
-sin, cos, and translate are the names of called functions.
-A rewrite rule can be applied whenever the pattern of function calls, literals, and variables on its left-hand side matches a code fragment.
-As shown, the two sides of a rewrite rule are specified as regular Julia code expressions, with the addition of unbound, optionally typed, variables.
-This syntax is able to cover all rewrites where a single expression tree is replaced with a new expression tree.
-This syntax does not allow rules matching multi-patterns (Jia et al., 2019; Yang et al., 2021).
-These patterns can match multiple expression trees and rewrite them each, potentially reusing variables in multiple output expressions.
-3.1. Incorporating Rewrites in the Compiler
+In many use cases, especially in program optimization, the extracted graph cannot contain cycles. In such cases, more refined extraction techniques are needed. Ours will be discussed in Section 3.5.
+Julia
+Julia is a general-purpose programming language that is primarily used for scientific computing.
+Being a dynamically typed language, functions can be defined for unconstrained argument types.
+A function can also be redefined for specific argument types (and counts) to express that it should behave differently for that combination of types.
+The different behaviors, i.e., the definitions of a function for different types and counts of arguments, are called methods in Julia.
+In essence, a function is an identifier, such as f or print, that it shares with all its methods.
+When a function call is to be executed, the Julia run-time system will look up and dispatch the function’s method that best matches the types and count of all the call’s run-time arguments.
+This is called multiple dispatch.
+In Julia, operators are syntactic sugar for functions.
+Even basic functionality like array indexing can be customized with methods for different types of arrays and indices.
+A method can, e.g., redefine how elements in a Symmetric matrix are accessed, avoiding the need to store symmetric elements twice.
+Methods that implement more abstract computations on matrices by accessing those matrices’ elements through index operations will automatically benefit from that custom access method when they are invoked on a Symmetric matrix.
+Because all operators are functions, Julia expressions correspond by and large to trees of function calls.
 On top, the Julia compiler pipeline is shown: Julia code is parsed into an AST, which is lowered to Julia IR, for which Julia Codegen produces LLVM IR, which is further compiled by LLVM Codegen.
 On the bottom, the figure zooms in on additional stages that are executed on the Julia IR.
-These are macro expansion and control flow destructuring to yield untyped IR, type inference and constant propagation that produce typed IR, inlining to obtain inlined IR and finally SROA and simplification to yield optimized IR.
-As in most compiler middle-ends, this optimization will operate on an intermediate representation (IR) that supports optimizations while still preserving high-level language-specific information (Lattner et al., 2020; Li et al., 2021; swiftlang, 2012).
+These are macro expansion and control flow destructuring to yield untyped IR, type inference and constant propagation that produce typed IR, inlining to obtain inlined IR and finally scalar replacement of aggregates (SROA) and simplification to yield optimized IR.
+Fig.
+2 shows Julia’s just-ahead-of-time compiler flow.
+At run time, when a method needs to be dispatched and executed, this flow is invoked on it.
+We will call this the current method.
+The compiler first lowers the method’s AST into an SSA-based IR.
+Because all operators are functions, that IR code consists mainly of control flow, phi-nodes, literals, assignments to SSA values, and function calls.
+It hence directly reflects all expressions from the source code, including all called functions and operators, the main difference being that each sub-expression result is assigned to its own SSA value in the IR.
+On this IR, the compiler applies type inference based on the run-time argument types, constant propagation, method inlining, and some other basic optimizations such as scalar replacement of aggregates (SROA).
+From this Julia IR, the compiler then generates LLVM IR, to which many more optimizations are applied by the LLVM compiler before it generates assembly code for the target architecture.
+Cranelift and Acyclic E-Graphs
+Cranelift (Fallin, 2023), a compiler back-end mostly used to generate code from Rust and WebAssembly, is the first production compiler to use e-graphs for code optimization.
+As Cranelift aims at JIT compilation, where compilation time is critical, they have introduced acyclic e-graphs (ægraphs).
+Like regular e-graphs, these represent multiple equivalent expressions in e-classes.
+But, while regular EqSat fully propagates the effects of introduced equalities, this is not the case with ægraphs.
+Equalities are greedily applied, and e-graph invariants are not fully restored.
+This means that some rewrites are potentially not discovered, but the compilation time is shorter and more predictable.
+More relevant to our work is the introduction of the CFG skeleton, a data structure kept alongside the e-graph to keep track of control flow instructions and instructions with side effects, two constructs that do not fit the EqSat paradigm for optimizing expressions.
+Its workings and how we leverage this technique in our work will be explained in Section 3.1.
+Because e-graphs in Cranelift are used in the compiler back-end, they model expressions extracted from the IR after inlining, which can be relatively large, but mainly contain simple, low-level instructions, many of which are pure.
+Cranelifts’ e-graphs and its use of the CFG skeleton are hence designed to conserve all control flow and side effects, instead focusing on the rewriting of pure expressions in between control flow and operations with side effects.
+3. A Rewrite-Based Compiler Middle-End
+Our system allows programmers to specify custom rewrite rules in Julia syntax.
+These rules are then applied on an e-graph representing the code of a method.
+Once the rules are applied and the e-graph is saturated or a timeout has been reached, the compiler will try to extract the optimal code embedded within that e-graph.
+Rules have a left-hand side and a right-hand side.
+Each side consists of a regular Julia expression, i.e., of a tree of function calls, plus variables and/or literals.
+Two example rules are and .
+In these rules sin, cos, ^, +, and translate are the names of the called functions.
+The ~ prefix identifies pattern variables in a rule.
+These can match any input values of the expression’s function calls, be it the values of other expressions, variables, or literals.
+The :: notation is used to specify the types of values for which the rule can be applied.
+If no type is specified, values of any type can be matched.
+A rewrite rule can be applied whenever the pattern of function calls, literals, and variables on its left-hand side matches an IR code fragment.
+As expressions in Julia source code and the SSA IR produced for them contain the same function calls, it is straightforward to search for such matches in a method’s SSA IR code.
+In other words, there is no need to have rule authors express equivalences in terms of IR constructs.
+They can instead express them in regular source code expressions similarly to how those would occur in source code to be optimized.
+This syntax using regular Julia code expressions, with the addition of unbound, optionally typed variables, covers all rewrites where a single expression tree is replaced with a new expression tree.
+This syntax does not allow rules matching multi-patterns (Jia et al., 2019; Yang et al., 2021).
+Such patterns can match multiple expression trees and rewrite them each, potentially reusing variables in multiple output expressions.
+As in many existing compiler middle-ends, the optimization based on these rewrite rules/patterns will operate on an intermediate representation (IR) that supports optimizations while still preserving high-level language-specific information (Lattner et al., 2021; Li et al., 2021; swiftlang, 2012).
 In our case, this is the Julia compiler’s SSA-based IR (Bezanson et al., 2012).
-Figure 2 shows the flow of the compiler.
-When a function needs to be executed, the just-ahead-of-time flow is invoked on it.
-We will call this the current function.
-The compiler first lowers the current function’s AST into an SSA-based IR.
-On this IR, it applies type inference, constant propagation, function inlining, and a small number of other basic optimizations.
-From this Julia IR, the compiler then generates LLVM IR, to which many more optimizations are applied by the LLVM compiler before it generates assembly code for the target architecture.
-There are two places where it makes sense to apply our rewrite optimizations: right before inlining, and after inlining and optimizations.
-Before inlining, the current function’s IR still contains function calls as they were written by the programmer.
-At this stage in the compilation, high-level rewrite rules can typically be matched and applied, i.e., rules involving functions of high-level abstract APIs.
-In some cases, it might be useful to also apply rewrites to code in which function calls have already been inlined.
-After inlining, it is possible that code from different function bodies down the call chain of the current function ends up in its IR.
-This can uncover additional rewrite opportunities.
-Our current prototype implementation performs the rewriting optimization both before and after inlining, but we note that the most interesting rewrites fire before inlining, when high-level API usage has not yet vanished from the code.
-For this reason, the example snippets of IR in this paper contain code that has not been inlined yet.
+A consideration to be made for pattern rewriting is at what point in the compilation to apply patterns.
+Concretely, patterns can be matched on IR before or after function calls have been inlined.
+Before inlining, the current method’s IR still contains function calls as they were written by the programmer.
+At that stage in the compilation, high-level rewrite rules can typically be matched and applied, i.e., rules involving functions of high-level, abstract, and domain-specific APIs.
+However, some rewrite opportunities could then remain hidden behind call barriers.
+This can happen, e.g., when APIs with domain-specific function names wrap more generic computations such as underlying linear algebra computations.
+If the programmer used the domain-specific wrapper APIs while the rewrite patterns are expressed in terms of the underlying linear algebra operations, those patterns will not be found in the IR code before inlining.
+Ideally, when exploring rewrite opportunities, the compiler should hence consider all possible combinations of inlined and not-yet-inlined functions.
+Support for such an exploration is for future work.
+Instead, our current prototype implementation performs rewriting optimization only before inlining.
+We note that many interesting rewrites that are not covered by classical compiler optimization occur before inlining, when high-level API usage has not yet vanished from the code.
+For this reason also, the example snippets of IR in this paper contain code that has not been inlined yet.
 3.1. From IR to E-Graph and Back Again
-Top left: a simple Julia function that computes x to the power n and that prints a message to illustrate the impact of side effects.
-Top right: the corresponding Julia IR that is generated if the power function is generated for two integer arguments.
-It consists of four basic blocks, one of which includes two phi functions.
-Bottom left: the e-graph for the expressions in the function consisting of 11 e-classes, including classes for the phi functions and for the side effect.
-Bottom right: the CFG skeleton of the function, which looks like the CFG itself, and in which each node refers to certain e-classes in the e-graph where the control flow depends on expressions modeled by those e-classes.
-To apply rewrites on the Julia IR of a function, we first convert it to an e-graph.
-For this, we employ the methods introduced in the Cranelift compiler (Fallin, 2023).
-Each function call is added to an e-graph, while control flow statements (goto, goto if not, phi, and return) are added to an external data structure called the CFG skeleton.
-Function calls that have side effects are also added to the CFG skeleton.
-Each statement in the CFG skeleton refers to the e-nodes on which it depends.
-Figure 3 shows an example Julia function pow together with its corresponding IR, e-graph, and CFG skeleton.
-In addition to the control flow statements, the call to println has been stored in the CFG skeleton because calling that function produces side effects.
-In this representation, the phi nodes are leaf nodes.
-Their dependencies are not included in the e-graph.
-This implies (i) that the e-graph does not contain sufficient information to reconstruct all control flow of the original code, and (ii) that rewrites including phi-node dependencies are not supported.
-It is the CFG skeleton that provides the extra information needed to reconstruct valid IR.
+Top left: a simple Julia method for function pow that computes x to the power n and that prints a message to illustrate the impact of side effects.
+Top right: the corresponding Julia IR that is generated if the pow method is generated for two integer arguments.
+It consists of four basic blocks, one of which includes two phi-functions.
+Bottom left: the e-graph for the expressions in the method consisting of 11 e-classes, including classes for the phi-functions and for the side effect.
+Bottom right: the CFG skeleton of the method, which looks like the CFG itself, and in which each node refers to certain e-classes in the e-graph where the control flow depends on expressions modeled by those e-classes.
+To apply rewrites on the Julia IR of a method, we first convert it to an e-graph and a CFG skeleton.
+For this, we leverage methods first introduced in the Cranelift compiler (Fallin, 2023).
+Fig.
+3 shows an example Julia function pow together with its SSA IR, e-graph, and CFG skeleton.
+The structured control flow present in the original source code has been lowered into four basic blocks in the Julia IR, each terminated by an instruction that jumps to another block or returns.
+Each e-node in the e-graph corresponds to an SSA value in the IR.
+For example, the value %7 in the IR corresponds to the e-node labeled -.
+The dotted boxes around the e-nodes represent the different e-classes.
+Since no equalities have been registered yet in this e-graph, each e-class consists of exactly one e-node.
+In this representation, the phi-nodes are leaf nodes, i.e., their dependencies are not included in the e-graph.
+This is because an SSA phi function introduces a branch in an expression that cannot be represented in an e-graph directly.
+The lack of phi-node information implies (i) that rewrites including phi-node dependencies are not supported, and (ii) that the e-graph by itself does not contain sufficient information to reconstruct all control flow of the original code.
+For (i), an example of a useful rewrite that is not supported is to reduce a phi-node to a value if all possible values in the phi-node are equivalent.
+For (ii), it is the CFG skeleton that provides the extra information needed to reconstruct valid IR.
+To that extent, the CFG skeleton contains all operations responsible for control flow and all operations with side effects, including function calls that have side effects.
+All operations in the skeleton reference the e-classes they depend on.
+To visualize these references, we use the roman numerals in Fig.
+3.
+Each roman numeral labels an e-class of which the corresponding SSA value is used in an operation with side effects, or in a control flow statement.
+For example, since a call to println has side effects, it is stored in the third block of the CFG skeleton, referencing e-class IV.
+Cranelift’s SSA IR format, like MLIR (MLI, 2019) and Swift (swiftlang, 2012), uses basic block arguments.
+These offer an alternative to phi-nodes for representing control-dependent data flow in SSA IRs.
+Whereas phi-nodes represent such data flow within a basic block, basic block arguments represent such data flow at the basic block level.
+Even though Julia IR contains phi-nodes, we opt to use block arguments in the CFG skeleton because all information regarding out-of-block uses of expressions that need to be materialized within a block is then explicitly represented within that block itself.
+To illustrate, the references to e-classes I, II, V, and VI in the two goto #2 statements in the skeleton are basic block arguments.
+They are passed as arguments to control flow statements and can be references in the destination block, similarly to how functions have arguments.
+The relation between basic block arguments and phi-nodes in the original Julia IR is simple.
+For example, the two phi-nodes in block #2 in the Julia IR correspond to the two basic block arguments passed in the CFG skeleton by blocks #1 and #3.
+Note that while the e-graph in Fig.
+3 contains nodes labeled phi, these are purely labels and do not carry the information contained in SSA phi-functions.
+The phi-nodes in the e-graph can be interpreted as being the corresponding basic block arguments.
+It is the CFG skeleton that allows the conversion back to valid Julia IR.
+As long as each statement in the CFG skeleton is elaborated, meaning that that statement and all of its dependencies are materialized in the generated IR, the resulting program has the same control flow and side effects as the original program.
 The need to support control flow stems from a number of common code patterns of expressions that are split over multiple basic blocks for the sake of code conciseness, readability, and maintainability.
@@
 - Keyword Arguments (kwargs):
-It is quite common that part of a function depends on a kwarg, e.g., to perform extra checks or to modify the behavior.
-See, e.g., the use of the flipkernel kwarg in NNLib.jl.111https://github.com/FluxML/NNlib.jl/blob/81e6cd1843dbd2baf2b7b2efd00e942f4f040a2d/src/fold.jl#L186-L191.
+It is quite common that part of a method depends on a kwarg, e.g., to perform extra checks or to modify the behavior. See, e.g., the use of the flipkernel kwarg in NNLib.jl (NNl, 2025).
 - LAPACK-style wrappers with uplo/fmt/trans/… chars:
@@
 Such checks are then independent of the values calculated beforehand.
-For example, in the LinearAlgebra.jl222https://github.com/JuliaLang/LinearAlgebra.jl/blob/master/src/bidiag.jl#L1414-L1448.
-standard library, the costly operation blks = findall(...) is first computed, after which there is a branch based on uplo, followed by a reuse of blks.
-Another example is CUDA.jl’s CUSPARSE wrappers: Plenty of API calls (which are a prime use case for e-graph based rewriting) occur outside of fmt checks, others occur inside such checks.333https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusparse/generic.jl#L48-L70.
+For example, in the LinearAlgebra.jl standard library (Lin, 2025), the costly operation blks = findall(...) is first computed, after which there is a branch based on uplo, followed by a reuse of blks.
+Another example is CUDA.jl’s CUSPARSE wrappers: Plenty of API calls (which are a prime use case for e-graph-based rewriting) occur outside of fmt checks, others occur inside such checks (CUD, 2025a).
 - API Version Checks:
-Continuing on the subject of API calls, it is common that certain API uses are conditional on the version of the library.
-For example, CUDA.jl includes checks such as CUSPARSE.version() v"11.7.2".444https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusparse/generic.jl#L304-L335.
+Continuing the subject of API calls, it is common that certain API uses are conditional on the version of the library.
+For example, CUDA.jl includes checks such as CUSPARSE.version() v"11.7.2" (CUD, 2025b).
 - Low-level Math:
@@
 The absolute cost of operations outside the if is then often not that great, but when it comes to optimizing very high-performance code, it can make a difference.
-Some examples can be found in the standard float.jl and math.jl libraries.555https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/float.jl#L320-L328666https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/math.jl#L756-L766.
+Some examples can be found in the standard float.jl (flo, 2025) and math.jl (mat, 2025) libraries.
 - Imperative Type Checks:
-For readability, or to avoid too many methods, imperative type checks are also sometimes used in Julia instead of multiple dispatch.
-Several examples can be found in the standard libraries.777https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusolver/linalg.jl#L35-L48, https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/abstractarray.jl#L2485-L2497, https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/intfuncs.jl#L592-L606.
 All of these patterns involve divergent control flow, which our work can handle without problems.
@@
 Otherwise, the previously materialized SSA value is reused.
-Like MLIR and the Swift Intermediate Language (SIL) (swiftlang, 2012; MLIR, 2019), Cranelift’s SSA IR format uses basic block arguments, which are bound by the branch operations in a block’s predecessors, to represent control-dependent data flow.
-Julia’s SSA IR, by contrast, uses phi-nodes.
-Basic block arguments are more convenient for materializing statements in the e-graph because all information regarding out-of-block uses of expressions that need to be materialized within a block are then explicitly represented within that block itself.
-For this reason, our code representation differs from existing uses of e-graphs and CFG skeletons in that it contains both phi-nodes and basic block arguments.
-Take, for example, the two goto #2 statements in the Julia IR in Figure 3.
-Since the destination block (#2) contains two phi nodes, the goto statements in the CFG skeleton pass the respective values as arguments.
-The e-graph itself still stores phi node values, but these are now handled as basic block arguments.
-During conversion from IR to e-graph and back from e-graph to IR, a conversion between phi nodes and basic block arguments is therefore performed.
+Our implementation uses Metatheory.jl (Cheli, 2021b), a Julia EqSat package based on egg (Willsey et al., 2021).
+At its core, it provides an efficient implementation of the e-graph datastructure and EqSat algorithms, and a system that allows to register new types of e-nodes.
+In our work, we use this system to implement e-nodes that represent Julia IR statements.
+Metatheory.jl also offers macros to easily specify rewrite rules with a high-level, declarative syntax.
+We extended this rule syntax to support type annotations, as will be discussed in Section 3.2.
+Lastly, Metatheory.jl contains a generic extraction algorithm.
+However, this algorithm is unsuited for our problem, as extraction needs to take into account the information from the CFG skeleton as well.
+In Section 3.5 we will discuss our extraction approach.
 3.2. Rules and Types
-One of the defining features of the Julia programming language is multiple dispatch.
-This feature allows users to define a new method of a function for a specific combination of argument types.
-The fact that the semantics of a function call are not only decided by syntax but also by the type of their arguments necessitates a way to represent this information in rewrite rules as well.
+With multiple dispatch, the semantics of a function call are decided not only by syntax (i.e., the function name), but also by the types and counts of their arguments.
+This necessitates a way to represent this information in rewrite rules.
 For example, a user might want to rewrite matrix multiplications into a function call to some external BLAS library.
-This means that a call to Base.:*, a function in Julia’s base library, should be rewritten, but only if its arguments are matrices.
-Users can specify type constraints in the rewrite rules by associating a type with an unbound variable. For example, a rule for rewriting a matrix multiplication can be written as
+This means that a call to Base.:*, a base library function, should be rewritten, but only if its arguments are matrices.
+Users can specify type constraints in the rewrite rules by associating a type with an unbound variable.
+For example, a rule for rewriting a matrix multiplication can be written as
+E-class analyses (Willsey et al., 2021) offer a framework to associate and propagate lattice information in an e-graph.
 We designed a novel e-class analysis for tracking types in the e-graph.
 For each e-class, this analysis maintains the most specific type that is compatible with all the e-nodes in that class.
-During e-graph construction, this type is simply the same as the types of statements in the source Julia IR.
+During e-graph construction, this type is simply the same as the types of statements in the Julia IR.
 When two e-classes with a different associated type are merged, the resulting e-class is determined by the typejoin function.
@@
 Similarly, if the type associated with an e-class changes, we rerun type inference for all e-classes containing e-nodes that depend on the changed e-class.
+As an optimization, for a small amount of different types that are joined, we construct a Union type that keeps track of all individual possible types.
+When the union grows too large, we fall back on a single type that is their most specific joint supertype.
+For example, if a value of type Integer is introduced in an e-class with associated type Float64, the new associated type becomes Union{Integer, Float64}.
+If, throughout EqSat, this union grows further, the associated type could become Number.
+Tracking small unions of specific types instead of joining them immediately into their joint supertype enables additional rewrite rules bound to the more specific types.
+Rewrite rules that change the type of a value produced by an expression can cause problems.
+If the value is used by a function that has no method defined for the new type, no valid code can be produced anymore.
+Our current implementation detects this when the type information of parent nodes is recomputed.
+When type inference finds no method for arguments with the new type, an error is thrown.
+It is the responsibility of the rule author to either ensure that rules only introduce new types that are compatible with all the functions that depend on the value, or that new methods are defined to make it so.
+In future work, we will implement the necessary support to skip the application of rewrite rules that can result in the production of invalid code.
 3.3. Constants in the E-Graph
@@
 They match an expression and replace that expression with a new one, plugging in any matched variable.
-Modern equality saturation frameworks such as egg (Willsey et al., 2021) and Metatheory.jl (Cheli, 2021b), which was used in this work, also support dynamic rewrites.
-When dynamic rewrite rules are applied, the right-hand side is first executed (potentially using additional e-class analysis data) and the original code is rewritten into the result of that execution, rather than into the right-hand side itself.
-When literals are part of the e-graph, dynamic rewrites can be used to perform simple constant propagation.
-In normal Julia code, literals of primitive types as well as composite types with a known size can appear directly in the IR.
+Modern EqSat frameworks such as egg (Willsey et al., 2021) and Metatheory.jl (Cheli, 2021b), which is used in this work, also support dynamic rewrites.
+When dynamic rewrite rules are applied, the expression in the right-hand side is first evaluated (potentially using additional e-class analysis data) and the original code is rewritten into the result of that execution, rather than into the literal right-hand side expression itself.
+When literals are part of the e-graph, dynamic rewrites can be used to perform simple constant folding.
+In normal Julia code, literals of primitive types and composite types with known size can appear directly in the IR.
 The values of these literals are embedded in the e-graph.
 The e-classes containing the corresponding e-nodes of the literals have an inferred type associated with them, but compared to e-classes containing nonliteral e-nodes, the actual value of that type is available as well.
-We extend rewrite rule type annotations with a special Comptime{T} type that allows to match against the type of the actual values in the e-graph instead of the inferred type stored in the e-class.
-Using instead of to signify dynamic rewrite rules, a rule such as
-will match any addition between two integer literals and insert the result of this addition in the respective e-class.
+To support constant folding of literals, we extend rewrite rule type annotations with a special Comptime{T} type that only allows to match against literals of type T.
+Using instead of to signify dynamic rewrite rules, a rule such as will match any addition between two integer literals and insert the constant-folded result of this addition in the respective e-class.
 Dynamic rewrite rules also open opportunities for partial evaluation-like optimization (Muchnick, 1998; Futamura, 1999) beyond the constant propagation capabilities of the Julia compiler.
@@
 With dynamic rewrite rules, a programmer can override this default behavior.
-To achieve this, we extended the rewrite rule format with rules of the form
-This rule introduces an equivalence between the function’s second argument in its e-graph and the constant W.
+To achieve this, we extended the rewrite rule format with rules of the form This rule introduces an equivalence between the function’s second argument in its e-graph and the constant W.
 A constant array holding W’s contents is added as an equivalent e-node to that argument’s e-class in the e-graph, to which dynamic rewrite rules can then be applied.
@@
 3.4. Rewrites in the Presence of Side Effects
-We depend on the static effect analysis built in the Julia programming language to decide whether a function call has side effects.
-This analysis determines several program properties for every statement in the IR of a function.
-The program property effect_free signals whether a statement is free from externally semantically visible side effects.
-A limitation of Cranelift’s approach using the CFG skeleton is that there is no way to rewrite instructions with side effects.
-In Cranelift, this is less of a concern as the code that is rewritten is lower level and typically contains a large amount of simple, low-level instructions, many of which are pure.
-In high-level Julia code, however, the IR we work on generally is much smaller and consists only of a few calls to high-level functions.
-Because these functions have more complex behavior than simple, low-level instructions, they also often include behavior with side effects.
-Still, it should be possible for developers to rewrite function calls that have side effects.
-For example, a user might want to rewrite an expression containing a function that allocates and returns an array.
-To enable this, we introduce a technique called CFG skeleton relaxation.
-We allow users to create rewrite rules that match against functions with side effects.
-When such a rule matches an expression in the e-graph and the rewrite is applied by adding the equivalence to the e-graph, all function calls that are part of the matched rule are removed from the CFG skeleton.
-However, the ability to ignore the effect analysis and overwrite function calls with side effects is not without risks.
-Users who are unaware of certain side effects might unknowingly remove them and run into hard-to-debug bugs.
-We mitigate this problem somewhat by requiring an explicit prefix unsafe in front of the rules that are allowed to remove elements from the CFG skeleton.
-Although this raises the bar somewhat to specify and use rewrite rules that unintentionally break the intended semantics of a rewritten piece of code, the risk of using rules incorrectly is obviously not mitigated completely.
-Semantic rule verification is outside the scope of this work.
-A rewrite rule might be semantically correct in one domain but incorrect in another, such as reduced precision in deep learning, so correctness is up to the rule author.
-Bad rules can introduce bugs.
+As discussed in Section 3.1, the CFG skeleton is used to ensure that all statements with side effects, such as the println in the example of Fig.
+3, are materialized when the final optimized program is extracted.
+We leverage Julia’s built-in static effect analysis to determine whether a function call has side effects.
+This analysis determines several properties for each IR statement.
+The program property effect_free signals whether a statement is “free from externally semantically visible side effects” (jul, 2025).
+For our purpose, however, this analysis is often too strict and too imprecise, resulting in some functions not being labeled effect_free, even though they do not have side effects that require them to remain in a program.
+For example, many functions that allocate memory are not marked effect-free, even if the allocated memory never gets accessed again.
+As a result, even if rewrite rules subsequently introduce equalities that would make calls to these functions superfluous, they will still be included in the optimized code.
+In short, there is a mismatch between Julia’s effect analysis and the desired freedom to operate of our rewrite system.
+In Cranelift, side effects are less of a concern because the code it rewrites is lower level and typically contains a large amount of simple, low-level instructions, many of which are pure.
+In high-level Julia code, by contrast, the IR we work on generally is much smaller and consists only of a few calls to high-level functions/methods.
+Because these functions/methods have more complex behavior than simple, low-level instructions, they also often include behavior with potential side effects according to Julia’s built-in analysis.
+To overcome this mismatch when rewriting Julia code, we allow users to prefix rewrite rules with the relaxed keyword.
+Whenever a relaxed rule matches code, all the function calls that are part of the CFG skeleton and that occur in the left-hand side of the rule are “detached” from the CFG skeleton.
+When the final optimized code is extracted from the e-graph, detached statements are not forced to be materialized.
+We call this technique CFG skeleton relaxation.
+Technically, it is possible to write relaxed rules that get rid of side effects that are depended upon by other parts of the code, thereby breaking its intended semantics.
+We note, however, that all practical rewrites that we encountered and that needed relaxation as we will discuss in Section 4, did not introduce problems.
+Finally, we note that semantic rule verification is outside the scope of this work.
+As a rewrite rule might be correct in one domain but incorrect in another, such as reduced precision in deep learning, correctness is up to the rule author.
 3.5. Extraction
-Example function where equality saturation reveals opportunities for reuse.
+Example function where EqSat reveals opportunities for reuse.
 Based on some condition, the function returns either the trace of matrix A or its determinant.
@@
 When Julia IR is generated from an optimized e-graph and its corresponding CFG skeleton, each statement in the CFG skeleton is considered a root node for e-graph extraction.
-This means that, in contrast to most work on e-graph-based optimizers, we essentially need to run multiple extractions to recover our program.
+This means that, in contrast to vanilla egraph extraction, we essentially need to run multiple extractions on the single e-graph to recover our program (Coward et al., 2023a).
 These extractions are not independent of each other.
-Take, for example, the code and corresponding IR and e-graph in Figure 4.
-The function f computes and returns two different values eig and x derived from matrix A.
+Take, for example, the code and corresponding IR and e-graph in Fig.
+4.
+The method f computes and returns two different values eig and x derived from matrix A.
 eig stores the eigen vectors and values, and x stores the trace or the determinant of the matrix, depending on control flow.
-In this example, two rules from a collection of linear algebra identities have been applied:
+In this example, two rules from a collection of linear algebra identities have been applied: and .
 These give rise to the extra e-nodes in the e-classes containing det and tr.
 Intuitively, we can see that an optimal extraction might be one in which the result of the call to eigen is reused to compute the determinant and/or trace instead of explicitly calling those functions.
-Indeed, that extraction is valid since the call to eigen dominates both the call to det and to tr so its result is available at those call sites.
+Indeed, that extraction is valid since the call to eigen dominates both the call to det and to tr, so its result is available at these call sites.
 We formulate the problem of finding an optimal extraction as an ILP problem.
 He et al.
-(He et al., 2017) propose an ILP description for e-graph extraction that ensures an acyclic extracted graph, which is important for further elaboration to produce linearized IR.
+(He et al., 2023) propose an ILP description for e-graph extraction that ensures an acyclic extracted graph, which is important for further elaboration to produce linearized IR.
 We adapt and extend their description to take into account e-node reuse from dominating extractions.
-An example E-graph with 5 e-classes.
-E-node n has 2 child e-classes.
-E-class c contains e-node n, and another, unnamed e-node.
-Both e-nodes in e-class c have the same two e-node parents.
-In this graph, nodes are additionally tagged with annotations that illustrate for e-node n the sets p(n) and chi(n).
 Let the e-graph be represented by a set of e-nodes , and a set of e-classes .
-Let be the set of e-nodes that are contained in e-class , and let be the set of children e-classes of a particular e-node .
-We associate with each e-node a set of e-nodes defined as:
-That is, the set of e-nodes that have a child e-class containing . Figure 5 shows an e-graph in which the different parts have been annotated with this notation.
+Let be the set of e-nodes contained in e-class , and let be the set of children e-classes of a particular e-node .
+We associate with each e-class a set of e-nodes defined as .
+That is, the set of parent e-nodes that have a child e-class .
+Fig.
+5 shows an e-graph and CFG skeleton in which the different parts have been annotated with this notation.
 We represent the statements contained in the CFG skeleton as the set .
 All of these statements cause a different expression to be extracted from the e-graph, rooted at each .
-For each statement the statements that dominate are represented by the set .
+For each statement , the statements that dominate are represented by the set .
 Each statement depends on a number of e-class arguments represented by the set .
-Finally, we associate a cost with every e-node such that our optimization problem becomes:888Changes to the formulation by He et al. (He et al., 2017) are highlighted in blue.
+Finally, we associate a cost with every e-node such that our optimization problem becomes111Changes to the formulation by He et al. (He et al., 2023) are highlighted in blue.
 Where, for each e-node and statement from the CFG skeleton, we have introduced a binary variable .
@@
 Furthermore, following the work by He et al.
-(He et al., 2017), each is a class cycle, a collection of classes that contain e-nodes that form a cycle by depending on other classes in the cycle.
+(He et al., 2023), each is a class cycle, a collection of classes that contain e-nodes that form a cycle by depending on other classes in the cycle.
 The e-nodes in an e-class that participate in a class cycle are represented by the set .
-For each class in each class cycle , we introduce the binary variables that are used to enforce acyclicity in the extracted graph.
+For each class in each class cycle , we introduce binary variables to enforce acyclicity in the extracted graph.
 is assigned 1 if none of the e-nodes in e-class that participate in the class cycle are selected.
-For a more complete overview and proof of acyclic extraction we refer to the work by He et al.
-(He et al., 2017).
+For a more complete overview and proof of acyclic extraction, we refer to the work by He et al.
+(He et al., 2023).
 To take into account the dominance information from the CFG skeleton, we adapted their formulation in a few ways.
@@
 Compared to the original ILP formulation by He et al.
-(He et al., 2017), we added an extra term to account for the fact that an e-node might have already been selected in a dominating extraction.
-The remaining constraints are completely analogous to the constraints by He et al.
-(He et al., 2017) except for the addition that each constraint is added separately for each possible extraction root .
-We implement this ILP formulation in JuMP (Lubin et al., 2023), a domain-specific language embedded in Julia for expressing and solving mathematical optimization problems.
-For a solver, we resort to HiGHS (Huangfu and Hall, 2018).
+(He et al., 2023), we added an additional term to account for the fact that an e-node might already have been selected in a dominating extraction.
+The other constraints are completely analogous to the constraints by He et al. (He et al., 2023) except for the addition that each constraint is added separately for each possible extraction root .
+We implemented this ILP formulation in JuMP (Lubin et al., 2023), a domain-specific language embedded in Julia to express and solve mathematical optimization problems.
+For a solver, we resorted to HiGHS (Huangfu and Hall, 2018).
 By default, our cost function currently defaults to 1, reducing the optimization problem to choosing the least number of e-nodes.
@@
 4. Use Case Demonstration
-To demonstrate the capabilities of our work, this section presents a number of use cases.
-This demonstration focuses on our main goal of allowing users and domain experts to express and exploit optimization opportunities in a convenient, flexible way.
-In other words, our use cases aim to demonstrate that our work can give developers the necessary control over optimizations without them having to become compiler experts and without restricting the flexibility and productivity they have come to expect from the Julia ecosystem.
+Example of domain-specific code used to apply 2D transformations to a point.
+The method performs a series of transformations: a horizontal translation over distance dx, a rotation over pi radians, a translation over cosine squared plus sine squared minus one, which equals zero, another rotation over pi radians, and a translation over minus dx.
+To demonstrate the capabilities of our work, this section presents a number of use cases, focusing on our main goal of allowing users and domain experts to express and exploit optimization opportunities in a convenient, flexible way.
+In other words, our use cases aim to demonstrate that our work can give developers the necessary control over optimizations without having to become compiler experts and without restricting the flexibility and productivity they have come to expect from the Julia ecosystem.
 In contrast, our work aims to give developers more flexibility, meaning less restrictions, to write code without incurring a reduction in optimization opportunities.
-Our work does not at all aim to squeeze more performance out of the Julia programming language and its compiler and run-time implementation.
-Indeed, any optimization that can be obtained with a rewrite rule can also be obtained by manually rewriting source code, such as by duplicating source code and then special-casing it.
-This section will hence not evaluate how much the performance of a piece of code can be improved with rewrite rules.
 4.1. Domain Specific Rewrites
-Example of domain-specific code used to apply 2D transformations to a point.
-The function performs a series of transformations: a horizontal translation over distance dx, a rotation over pi radians, a translation over cosinus squared plus sinus squared minus one, which equals zero, another rotation over pi radians, and a translation over minus dx.
-With a handful of rules, these can be reduced to the identity transformation.
-Figure 6 shows a function that applies a sequence of similarity transformations on a 2D point.
-On close inspection, application of geometric properties, trigonometric identities, and algebraic simplifications can lead to the reduction of all these transformations to a single identity transform.
-Indeed, the translation on line 8 can be simplified to p by applying
+Fig.
+6 shows a method that applies a sequence of similarity transformations on a 2D point.
+On close inspection, under the assumption that exact floating point results must not be maintained, application of geometric properties, trigonometric identities, and algebraic simplifications can lead to the reduction of all these transformations to a single identity transform.
+Indeed, the translation on line 8 can be simplified to p by applying the following three rules: , , and .
 After this translation has been eliminated, the two rotations by radians similarly end up not transforming the point.
 Lastly, the first and last rotations now cancel each other out as well.
-After all the corresponding rewrites, the optimized function simply returns its argument p.
-With this example, we also illustrate the necessity of CFG skeleton relaxation.
-The effect inference in the Julia compiler is unable to reason through the rotation matrix construction in the rotate function on line 2 in Figure 6.
-Because of this, the rewrite rule author is required to allow CFG skeleton relaxation for those rules where rotate is called by prefixing those rules with unsafe.
+After all the corresponding rewrites, the optimized method simply returns its argument p.
+This example also illustrates the necessity of CFG skeleton relaxation.
+The Julia compiler’s effect inference fails to reason through the rotation matrix construction in the rotate function on line 2 in Fig.
+6.
+As a result, the allocation of the rotation matrix is conservatively considered to have side effects, which means that, by default, even if an optimization makes the rotation matrix obsolete, the matrix will still be materialized by the optimized code.
+The rewrite rule author is hence required to allow CFG skeleton relaxation for rules with rotate by prefixing those rules with relaxed.
 4.2. Aiding Multiple Dispatch
-Instead of using rewrites purely for simplifying code, they can also be used to generate code that provides additional information to the Julia compiler’ type system to trigger the execution of better performing function implementations through multiple dispatch.
+Instead of using rewrites purely for simplifying code, they can also be used to generate code that provides additional information to the Julia compiler’ type system to trigger the execution of better performing methods through multiple dispatch.
 An example in the domain of linear algebra is to automatically wrap certain matrix expressions in a new, more specific type that allows more efficient implementations for subsequent computations.
@@
 For the case where and are the same matrix, that is, , we know that the result is a symmetric matrix.
-We can encode this fact by applying the following rewrite rule:
+We can encode this fact by applying the following rewrite rule: .
 The Symmetric function from the LinearAlgebra.jl package takes the upper triangle of its argument and uses that to efficiently represent a symmetric matrix.
 More importantly, it returns a value of type Symmetric.
-Other functions in that linear algebra package, e.g., for solving eigenproblems, have specialized implementations for symmetric matrices.
-After rewriting with the above rule, those specialized implementations will now be called automatically thanks to the multiple dispatch system knowing that they are invoked on type Symmetric instead of on a more generic matrix type.
-Another example is the expression , which is symmetric if B is symmetric as well. Again, this fact can be encoded with the following rule:
-Both these examples build on the abstractions of the Julia standard library, but the same concepts can also be applied to operations and types in other, external packages.
+Other functions in that linear algebra package, e.g., for solving eigenproblems, have specialized methods for symmetric matrices.
+After rewriting with the above rule, those specialized methods will now be called automatically thanks to the multiple dispatch system knowing that they are invoked on the type Symmetric instead of on a more generic matrix type.
+Another example is the expression , which is symmetric if B is symmetric as well, as can be encoded with the rule .
+Both examples build on the abstractions of the Julia standard library, but the same concepts can also be applied to operations and types in other, external packages.
 4.3. Multi-Line Broadcast Fusion
-Julia code that uses broadcasting syntax: using the dot operator, a elementwise ReLu operation and an elementwise addition are applied to vectors A and B.
+Julia code that uses broadcasting syntax: using the dot operator, an elementwise ReLu operation and an elementwise addition are applied to vectors A and B.
 In other words, the relu function is broadcasted over matrix A and vector B is added to the result using broadcasting addition.
@@
 Julia’s broadcasting mechanism allows users to apply functions element-by-element on one or more arguments containing multiple elements.
-Syntactically, this is done by adding a .
-between the function and its argument list.
-The top of Figure 7 shows two broadcast operations.
+Syntactically, this is done by adding a dot (.) between the function and its argument list.
+The top of Fig.
+8 shows two broadcast operations.
 relu.(A) applies the scalar relu function to each element of the matrix A.
@@
 Although the two arguments to .+ have a different shape, the operation can take place because the shapes are compatible.
-Broadcasting in Julia is implemented by transforming expressions containing this special dot syntax into calls to the standard library functions broadcasted and materialize.
-Figure 8 shows the result of this transformation on the broadcasting expression in Figure 7.
+The Julia compiler transforms expressions containing this special dot syntax into calls to the standard library functions broadcasted and materialize.
+Fig.
+8 shows the result for the expression in Fig.
+8.
 A call to broadcasted builds a lazy representation of the broadcast result.
@@
 Rewrite rules make this easy.
-Take for example the rule
-This rule will remove the intermediary materialization for any call of the form f.(x) where x is a variable that was defined on a different line with a broadcast expression itself.
-In the example code in Figure 9, fusion will now be applied despite the two dot operators occurring in two different statements on two separate lines.
-This example illustrates how the system we propose not only supports domain-specific optimization, but at the same time enables optimization of common code patterns involving only core Julia primitives.
-An important detail is that here, once more, CFG skeleton relaxation is required, since materialize ends up calling a foreign C function function which the compiler cannot deem effect-free, poisoning the remainder of the effect analysis.
-In practice, this means the rule needs to be prefixed with unsafe.
+Take, for example, the rule This rule will remove the intermediary materialization for any call of the form f.(x) where x is a variable that was defined on a different line with a broadcast expression itself.
+In the example in Fig.
+10, fusion will now be applied despite the two dot operators occurring in two different statements on two separate lines.
+This illustrates how the system we propose not only supports domain-specific optimization but also at the same time enables optimization of common code patterns involving only core Julia primitives.
+An important detail is that here, once more, CFG skeleton relaxation is required, since materialize ends up calling a foreign C function which the compiler cannot deem effect-free, poisoning the remainder of the effect analysis.
+In practice, this means that the rule needs to be prefixed with relaxed.
 Using PkgEval.jl, a tool to automatically evaluate the tests of Julia packages, we evaluated how many times the above rule’s pattern occurs in real-life code in 3153 different packages.
@@
 For other patterns, e.g., an expression where the outer broadcasted function is a binary function, separate rules need to be written.
-Table 1 gives an overview of how many times different multi-line broadcasting patterns were found in our analysis.
-The three last patterns are different configurations of broadcasting a binary function.
+Table 1 lists how many times our analysis found different multi-line broadcasting patterns that can all be rewritten to broadcasted(~f, ~x) and broadcasted(~binop, ~x, ~y).
+The last three patterns are different configurations of broadcasting a binary function.
 Note that the matches for these three patterns are not independent: a match against the last pattern, which will result in two invocations of materialize being optimized out, implies the two other patterns can be matched and applied as well, each optimizing out only one materialize invocation.
@@
 For example, instead of writing deep learning layers and non-linear activation functions using core language constructs, programmers often resort to using frameworks or libraries that contain optimized implementations of matrix multiplications, activation functions, attention computation, and others (Elshawi et al., 2021; Parvat et al., 2017).
-This works well for most common use cases but falls apart when a programmer requires custom behavior that is not yet supported by the framework.
+This works well for most common use cases, but falls apart when a programmer requires custom behavior that is not yet supported by the framework.
 In those cases, the programmer is required to rewrite parts of the program to stop using the framework and instead depend on a different framework or the core language.
-Instead of having to opt out of framework functionality and rewriting code, it would be much better if one could write simple, readable code and have the frameworks themselves automatically discover opportunities to use their better performing implementations.
-Our rewrite system gives framework authors the option for implementing such discovery and optimization almost for free.
-As an example, take the code in Figure 9 that shows an implementation of the forward pass of a single fully connected feedforward layer.
-There already exist several Julia packages that allow us to compute such operation more efficiently (Faingnaert et al., 2022; Innes et al., 2018; Innes, 2018).
-For example, the programmer could rewrite the code as shown in Figure10 to use the optimized fused_dense_bias_activation function from the Lux.jl deep learning framework (Pal, 2023).
+Instead of having to opt out of framework functionality and rewriting code, one could ideally write simple, readable code and have the frameworks themselves automatically discover opportunities to use their more performant implementations.
+Our system gives framework authors the option for implementing such discovery and optimization almost for free.
+As an example, the code in Fig.
+10 shows an implementation of the forward pass of a single fully connected feedforward layer.
+Several existing Julia packages allow us to compute such operations more efficiently (Faingnaert et al., 2022; Innes et al., 2018; Innes, 2018).
+For example, the programmer could rewrite the code as shown in Fig.10 to use the optimized fused_dense_bias_activation function from the Lux.jl deep learning framework (Pal, 2023).
 A disadvantage of such manual rewriting is that the code becomes less readable for someone not familiar with the Lux.jl APIs.
 Moreover, the use of that specific framework’s optimized implementation, which is in essence a decision about a non-functional aspect of the code, is then hard-coded in the code that specifies the functionality.
-This requires the developers and the maintainers of this code to be knowledgable in the code’s application domain as well as in the Lux.jl framework.
+This requires the developers and the maintainers of this code to be knowledgeable in the code’s application domain as well as in the Lux.jl framework.
 Switching to other frameworks in the future will require rewriting code.
 Clearly, hard coding the dependence on Lux.jl and its APIs in the code has a number of downsides.
-With our system, instead of rewriting the original code of Figure 9, it suffices to add a rewrite rule that maps the original code pattern onto the optimized fused_dense_bias_activation function:999Type annotations have been left out for the sake of conciseness of this rule, but they are needed to make sure to only match when W and x are matrices, and b is a vector.
-| materialize(broadcasted(~, broadcasted(+, ~W * ~x, ~b))) | ||
-Combined with the multi-line broadcast rewriting rule explained above, which rewrites the original code into a single line without the intermediary variable logits, this rule maps the single-line broadcast expression onto a call to the library function LuxLib.fused_dense_bias_activation, which will yield compiled code equivalent to what would be obtained with the code in Figure 10.
+With our system, instead of rewriting the original code of Fig. 10, it suffices to add a rewrite rule that maps the original code pattern onto the optimized fused_dense_bias_activation function:
+| materialize(broadcasted(~, broadcasted(+, ~W::AbstractMatrix * ~x::AbstractMatrix, ~b::AbstractVector))) | ||
+Combined with the multi-line broadcast rewriting rule explained above, which rewrites the original code into a single line without the intermediary variable logits, this rule maps the single-line broadcast expression onto a call to the library function LuxLib.fused_dense_bias_activation, identical to what would be generated for the code in Fig.
+10.
 This example illustrates how our approach allows for a better separation of concerns.
 Another advantage is that alternative rewrite rules can be provided for different Julia packages that provide different optimized implementations for similar functionality.
-Different rewrite rules also don’t obfuscate each other.
+Different rewrite rules also do not obfuscate each other.
 All rewrite rules can be represented in the e-graph.
@@
 As such, rewrite rules from different authors are fully interoperable.
-Similarly, without changing the source code, different rewrites can still occur by applying different sets of rewrite rules that can for example be offered by different Julia packages.
-This example also illustrates how our system supports the composition of high-level, domain-specific optimizations with optimizations of core Julia primitives.
-In this way, our system’s scope is much broader than recent equality saturation based optimizers such as Cranelift.
+Similarly, without changing the source code, different rewrites can still occur by applying different sets of rewrite rules that can, for example, be offered by different Julia packages.
+This example also shows how our system supports composing high-level, domain-specific optimizations with optimizations of core Julia primitives.
+In this way, its scope is much broader than recent EqSat optimizers like Cranelift.
+4.5. Keeping up with evolving and emerging libraries for GEMMs
+As an example of multiple libraries providing complementary functionality, consider General Matrix Multiplications (GEMMs), a cornerstone in many scientific domains.
+GEMMs are also often combined with element-wise operations or elops on their operands or result, for example with non-linear activation functions such as ReLU.
+Many libraries exist to perform different GEMM variants (Van Zee and van de Geijn, 2015; Faingnaert et al., 2022; Torres et al., 2024), each offering different trade-offs and interfaces.
+For targeting NVIDIA GPUs, e.g., the cuBLAS library provides the highest performance for basic GEMM expressions on matrices with simple datatypes.
+It also offers some kernels that fuse often-used elops with the GEMM.
+For elops not supported by cuBLAS, the alternative GemmKernels.jl can be used (Faingnaert et al., 2022).
+This package leverages Julia’s multiple dispatch and GPU programming support (Besard et al., 2019) to generate performant GEMM code on the fly, including for user-defined elops and for GEMMs on less commonly used data types such as dual numbers, which are useful for automatic differentiation (Revels et al., 2016).
+Over time, such libraries evolve.
+For example, cuBLAS has grown to include ever more kernels with fused elops.
+Ideally, programmers should not have to rewrite their application code to make use of newly available functionality or when switching from one library to another.
+Instead, they should be able to write generic code that automatically makes use of the available functionality in libraries.
+Rewrite rules enable this.
+Take for example the expression C .= relu.( * (f_a.(A) * f_b.(B)) + * C), where A, B, and C are matrices and and are scalars.
+The different GEMM and linear algebra packages can each provide a custom set of rewrite rules that rewrite as much as possible of that expression into calls to optimized functions.
+For example, a rule added to the LinearAlgebra.jl package would allow this expression to be rewritten into a more efficient call to mul!
+that carries out the matrix multiplication, scaling by and , and accumulation into C at once.
+Then only the elops (f_a, f_b, and relu) remain to be executed in separate kernels.
+For particular argument types, such as 16-bit floating-point matrices in GPU memory, rewrite rules in the GemmKernels.jl package would even allow the complete expression to be fused into one single kernel.
+Other packages could still contain rewrite rules for other GEMM configurations or variants.
+Depending on which rules are loaded, the same code hence can be rewritten in different ways, exploiting the various and possibly evolving optimization opportunities that complementary libraries offer without having to rewrite the source code.
 5. Compilation Time Evaluation
 Julia is a just-ahead-of-time compiled language for scientific computing.
-By default, Julia automatically compiles a native, type-specialized version of a function just ahead of the first time it is to be executed on some combination of argument types.
-Julia hence positions itself somewhere between typical ahead-of-time compilers such as LLVM and gcc and just-in-time compilers such as Cranelift or the compilers used in a Java VM.
+By default, Julia automatically compiles a native, type-specialized version of a method just ahead of the first time it is to be executed on some combination of argument types.
+Julia hence positions itself somewhere between typical ahead-of-time compilers such as LLVM and gcc and just-in-time compilers such as Cranelift or those used in a Java VM.
 Compilation speed is hence important for Julia.
-Alternatively, a user can request ahead-of-time compilation of functions and optionally invoke our rewriting rule-based optimization, depending on whether or not the investment in compilation time might yield a positive return on investment.
-As scientific software most often has relatively long execution times, compilation times are therefore not as critical as in more traditional JIT compilers.
+A user can also request ahead-of-time compilation of methods and optionally invoke our rewriting rule-based optimization, based on whether or not the investment in compilation time might yield a positive return on investment.
+As scientific software often has relatively long execution times, compilation times are hence not as critical as in traditional JIT compilers.
 Still, we aim for interactive compilation speeds with our work to allow for rapid prototyping.
-5.1. Complexity of E-Graph Construction
-In general, termination for equality saturation is not guaranteed since application of rewrite rules can introduce new rewrite opportunities indefinitely.
-(Willsey et al., 2021; Suciu et al., 2025) In practice, this issue is most often tackled by equality saturation timeouts in combination with heuristic schedulers that ensure that all rewrite rules get a chance to be applied by implementing a rule application back-off mechanism (Cheli, 2021a; Yang et al., 2021).
-The non-terminating behavior of equality saturation is due to repeated application of rule application and rebuilding the e-graph by joining newly discovered equivalences.
-The actual construction of the e-graph, however, is not concerned with this loop and does, in fact, execute in polynomial time.
-We reimplemented the algorithm introduced for Cranelift (Fallin, 2023) in Julia.
+5.1. Complexity Analysis of Equality Saturation
+E-Graph Construction
+We reimplemented the algorithm that converts IR into an e-graph introduced for Cranelift (Fallin, 2023) in Julia.
 This algorithm starts by computing the dominator tree of the input SSA code.
@@
 At the same time that the e-graph is being constructed, the CFG skeleton is being built.
-This involves pushing the corresponding e-classes of effectful statements in a CFG structure which has a worst case runtime complexity of , where is the code size, but an amortized complexity of .
-Since no equalities are introduced yet during e-graph construction time, inserting new e-classes in the e-graph does not require running the relatively expensive rebuilding procedure.
-As such, e-graph construction only requires inexpensive hash map lookups and insertions for each statement, which have a worst case runtime complexity of but an expected worst case runtime of .
-Overall, the worst case runtime complexity of our e-graph construction is but on average and in practice, as will be discussed in Section 5.2, the worst case runtime complexity is .
+This involves pushing the corresponding e-classes of effectful statements in a CFG structure which has a worst-case execution time complexity of , where is the code size, but an amortized complexity of .
+Since no equalities are introduced yet during conversion from IR to e-graph inserting new e-classes in the e-graph does not require running the relatively expensive e-graph rebuilding procedure.
+As such, e-graph construction only requires inexpensive hash map lookups and insertions for each statement, which have a worst-case execution time complexity of but an expected worst-case execution time of .
+Overall, the worst-case execution time complexity of our e-graph construction is but on average and in practice, as will be discussed in Section 5.2, the worst-case execution time complexity is .
+E-Graph Saturation
+In general, termination for EqSat is not guaranteed because the application of rewrite rules can introduce new rewrite opportunities indefinitely (Willsey et al., 2021; Suciu et al., 2025).
+In practice, this issue is most often addressed by EqSat timeouts in combination with heuristic schedulers that ensure that all rewrite rules get a chance to be applied by means of a rule application back-off mechanism (Cheli, 2021a; Yang et al., 2021).
+The non-terminating behavior of EqSat is due to repeated rule application and rebuilding the e-graph by joining newly discovered equivalences.
+Extraction
+Constructing the ILP problem entails a few different steps.
+First, simple cycles are determined using Johnson’s algorithm (Johnson, 1975), and the dominator tree is constructed.
+Since the number of cycles in a graph can grow exponentially in the number of nodes (Johnson, 1975), in the worst case, the time required to construct the problem is exponential as well.
+In practice, however, e-graphs typically contain only a few short simple cycles.
+Next, the minimization objective itself can be constructed in , with the number of e-nodes, and the number of statements in the CFG skeleton.
+Assuming a low number of class cycles, the problem construction time is asymptotically dominated by the construction of the class constraint (Equation 3).
+In the worst case, this takes time, with the maximum number of children e-classes of an e-node, the maximum number of dominating statements, and the maximum number of e-nodes in an e-class.
+Solving the ILP problem is an NP-hard problem, but as will be shown in Section 5, solving times are of the same order of magnitude as construction.
 5.2. Experimental Compilation Time Evaluation
-A graph showing the compilation time for different parts of our rewrite optimizations.
-The graph consists of three vertical bar charts next to each other, representing time spent in the pre-inlining, post-inlining, and finalizations stage.
-Each bar chart is divided in different colors indicating the time needed for specific compilation steps.
-The different compilation steps are: IR to e-graph conversion, e-graph saturation, e-graph to ir conversion, type reinference, invocation queueing, ILP problem construction, ILP solver (HiGHS), optimized function insertion, inlining pass, llvm optimization and code generation, and other.
-Figure 11 shows the compilation times spent in different steps of the whole optimization pipeline for optimizing the forward function from Figure 9.
-Important to note is that these numbers include the time spent optimizing each function in the call graph of the top-level function.
-For the forward method, the call graph consists of 43 methods that each pass through the full optimization pass.
-The chart partitions the optimization flow and its compilation time into three stages.
-All benchmarks were run on a machine with a 6-core Intel i7-9750H processor and 24GB RAM.
-The code was run on Julia v1.11.0 (commit 501a4f25c2b), with LLVM 16.0.6.
-5.2.1. Pre-Inlining
-In this first stage, rewrites are applied before the code body of the function is inlined.
-Due to the small number of rules, e-graph saturation only takes a small portion of the work in this stage.
-Similarly, ILP problem construction and solving runs quickly because non-inlined code consists of few statements, leading to a small ILP problem to solve.
-In this stage, a larger part of the run time is spent reinferring the types in the generated IR.
-This also includes replacing dynamic function calls with static invocations of the correct methods if enough type information is available to make this decision.
-The last step of this stage consists of collecting the non-inlined IR of each method that is invoked, and adding this IR to the top of a worklist.
-Methods invoked within a function end up at the top of the worklist and are fully optimized before the second stage of the current function can begin.
-5.2.2. Post-Inlining
-The second stage starts by replacing each method invocation with the optimized variant if rewrites were applied to it.
-Next, we have the default inlining pass that is also invoked during regular Julia compilation.
-In this second stage, conversion from IR to e-graph takes much longer because the inlined code consists of many more nodes.
-For the same reason, e-graph saturation takes longer as well.
-Because there are no rules to apply post-inlining in this instance, no time is spent on ILP, and type re-inference runs faster.
-To assess the scalability of the conversion from IR to e-graph and back, Figure 12 shows those times for IRs of different sizes.
-The IR dataset was collected by running the optimizer on the factorize function from Julia’s linear algebra standard library and on the many functions that are part of this function’s call graph.
-Both conversions clearly scale linearly with the number of statements.
-E-graph to IR conversion does not meaningfully show up in the compilation times in Figure 11 because this operation is only executed if actual changes to the e-graph were made, which happens infrequently in this instance.
-Similarly, re-inference times in the post-inlining stage do not take much time because this step is only applied when the e-graph has changed.
-This is in contrast to the pre-inlining stage, where part of the re-inference procedure needs to be executed either way to prepare the IR for inlining in the next stage.
-A scatter plot showing the time needed to convert from IR to e-graph and vice versa.
-The x-axis shows the number of IR statements, the y-axis shows the time needed in microseconds.
-Datapoints for IR to e-graph conversion are shown with an orange, plus-shaped marker, and e-graph to IR is shown using an red, triangle-shaped marker.
-Both cases a similar linear trend upwards.
-5.2.3. Finalization
-During the last stage, the functions that were enqueued in the previous stage have been optimized and are inserted if rewrites were applied to them.
-Finally, compilation is passed on to LLVM to generate an executable, native function from the optimized IR.
-This step takes more than a third of the total time.
-In Figure 11, Invocation Queueing, Inlining, and LLVM Compilation are steps that are not specific to our compilation flow but are also taken in the native Julia compiler.
-In the example discussed above, 80% of the total time is spent in these steps.
-The Julia compiler hides much of the compilation costs by caching the compiled code.
-Tighter integration of our optimization into the Julia compiler could thus lead to faster compile times as well.
-As briefly noted in Section 3.1, rewrite rules will be applied more often on IR before inlining because that IR contains function calls roughly as they were written by a programmer.
-In IR to which inlining has been applied, on the other hand, it is difficult to anticipate which function calls will end up visible, making it less productive to write rules targeting IR at this level.
-For this reason, a worthwhile trade-off might be to disable rewrites in the second, post-inling stage.
-This would eliminate the high cost of the IR to e-graph conversion that is included in the middle bar in Figure 11.
-5.2.4. ILP Construction and Solving
-Table 2 shows the time required to construct and solve the ILP problem for several examples discussed in this paper.
-With the exception of the code for simplifying 2D transformations in Figure 6, e-classes do not contain many equivalences, leading to fast problem construction and solving.
+We conducted our measurements on a 64-core AMD Ryzen Threadripper PRO 7985WX and 502GiB RAM, with Julia Version 1.11.5 (LLVM 16.0.6).
+Fig.
+11 shows the time taken for the different steps of our optimization scheme for Fig.
+10’s forward function (invoked with training=false) and the rewrites to fuse the operations as discussed in Section 4.4.
+Conversion
+E-graph conversion, both from and to Julia IR, takes relatively little time for this example.
+To evaluate the run time of conversions for other code, Fig.
+12 shows the time needed to convert between IR and E-Graph for many different Julia methods.
+These methods were selected arbitrarily by collecting all the methods in the call graphs encountered during the compilation of some methods from Julia’s core and linear algebra standard library.
+While our optimization targets IR without inlined function calls, we also include timing results for IR with inlined calls to capture performance on longer IR.
+Fig.
+13 shows that the vast majority of the evaluated methods consists of only a handful of IR statements.
+This means that we typically do not expect large e-graphs during optimization.
+Code Generation
+Not shown in the compilation time overview is the time that LLVM spent to optimize and generate the final executable code from Julia IR.
+The reason for the omission in the figure is that this step also occurs for regular Julia code compilation.
+For the example from Fig.
+11, LLVM code generation takes around 75ms, or about 95% of total compilation time.
+For regular Julia code, much of the compilation cost imposed by LLVM is hidden by caching of compiled code.
+Tighter integration of our optimization into the Julia compiler could similarly help hide this cost as well.
+ILP Construction and Solving
+Fig.
+11 shows that most of the time is occupied by constructing and solving the ILP extraction problem presented in Section 3.5.
+Apart from the inherent computational complexity of solving the ILP problem, the conversion takes a large amount of time for converting the e-graph into a collection of variables and constraints in an external ILP solver library.
+Table 2 shows the time required to construct and solve the ILP problem for several use cases discussed in this paper.
+With the exception of the code for simplifying 2D transformations in Fig.
+6, e-classes do not contain many equivalences, leading to fast problem construction and solving.
 For the 2D transformation example, particular rewrite rules, e.g., those that combine two subsequent rotations into one rotation, give rise to an infinite number of rewrite opportunities.
@@
 In the future, better scheduling of rules and detection of these cases could limit the amount of e-graph blow-up as well.
+| Construction [ms] | Solve [ms] | Variables | Constraints | E-Classes | E-Nodes | |
+| Fig. 4 | 0.46 | 1.042 | 126 | 364 | 26 | 28 |
+| Fig. 6 | 24.16 | 208.66 | 918 | 5325 | 337 | 1,002 |
+| Fig. 10 | 0.37 | 5.02 | 48 | 168 | 21 | 23 |
+6. Performance Evaluation
+| Use Case | Description | Speed-up |
+| Fig. 6 | 2D transformations | |
+| Fig. 4 | Linear algebra value reuse | up to |
+| Fig. 10 | Fused neural network | up to |
+| Sec. 4.5 | GEMM rewrites | up to |
+Our work does not at all aim to squeeze more performance out of the Julia programming language and its compiler and run-time implementation.
+Indeed, any optimization that can be obtained with a rewrite rule can also be obtained by manually rewriting source code, such as by duplicating source code and then special-casing it.
+Still, as the goal of the rewrite rules is to enable optimizations, this section reports the performance gains that can be achieved for some of the previously presented use cases.
+Our results are summarized in Table 3.
+The use case of Fig.
+6 where domain-specific code representing a 2D transformation is optimized into an idempotent function shows a large speed-up, as expected by all computations being optimized out of the function.
+The other examples paint a more interesting picture.
+For the use case of Fig.
+4, where the computation of a matrix determinant is simplified by reusing the already computed eigenvalues, Fig.
+14 shows a moderate speed-up across the board.
+For problem sizes around , this speed-up is more pronounced.
+For reference, yellow dots show the speed-up when simply omitting the determinant computation completely.
+In other words, they show what speed-up can be achieved by reducing the execution time of the determinant computation to zero.
+The fact that those roughly match with the optimized code shows that in the rewritten code, the determinant value basically comes for free, whereas it costs up to 10% in the original code.
+Similarly, for the use case of Fig.
+10 where the forward pass of a neural network is replaced with a fused implementation, the plot in Fig.
+16 shows a speed-up for small problem sizes but ends up about as fast as the original code for larger problem sizes, as the time spent in the matrix multiplication tends to dominate.
+222For problems involving GPU computation (Fig 16, Fig.
+16), we measured execution time on a machine with a 16-core Intel Xeon E5-2637 v2, NVIDIA GeForce RTX 2080 Ti (12GB), and 64GB RAM, with Julia Version 1.11.5 (LLVM 16.0.6).
+For the other problems, we use the same setup as in Section 5.
+Finally, we look at the example of rewriting the expression C .= relu.( * (f_a.(A) * f_b.(B)) + * C), discussed in Section 4.5.
+Fig.
+16 shows that the basic linear algebra rewrite, which fuses everything except the elops (f_a, f_b, relu) into one GPU kernel.
+This provides a moderate speed-up across the board over the baseline implementation in which each operation in the whole expression corresponds to a separate GPU kernel.
+The GemmKernels rewrite, which fuses all operations into a single kernel, is worthwhile for smaller problem sizes, and shows a small slowdown for larger problems, where the core GEMM computation dominates.
+This performance evaluation of four uses cases demonstrates that our system can indeed be used to achieve significant speed-ups by exploiting domain-specific optimization opportunities, in a programmer-friendly manner.
 7. Limitations
 Control Flow and Inlining
-As discussed in Section 3.2, our approach to converting IR to an e-graph using the CFG skeleton currently does not allow applying rewrites that alter control flow.
-Similarly, rewrite opportunities where part of an expression is hidden behind a function call are in general not considered.
-We optimize each function separately, applying rewrite rules both before and after inlining is applied.
-Other intermediate representations, discussed further in Section 7, could be used to represent the entire program in an e-graph and allow rewriting across call boundaries and control flow.
+As discussed in Section 3, our approach to converting IR to an e-graph using the CFG skeleton currently does not support rewrites that alter control flow.
+Similarly, since we apply rewrites on code where function calls have not yet been inlined, rewrite opportunities where part of an expression is hidden behind a function call are not considered.
+We optimize each method separately.
+Other intermediate representations, discussed in Section 8, could be used to represent the entire program in an e-graph and allow rewriting across call boundaries and control flow.
 Saturation and Extraction
 In this work, we have proposed an ILP formulation for optimal e-graph extraction.
-The problem of extracting e-nodes from a general e-graph is NP-hard, which means that there is no guarantee that an optimal solution can be found (Goharshady et al., 2024).
-Currently, when the ILP solver is unable to find a solution, we fall back to a greedy extraction algorithm that is built into Metatheory.jl.
+The problem of extracting e-nodes from a general e-graph is NP-hard (Goharshady et al., 2024), which means that there is no guarantee that an optimal solution can be found in polynomial time.
+Currently, when the ILP solver is unable to find a solution within a given timeframe, we fall back to a greedy extraction algorithm that is built into Metatheory.jl.
 In this work, we chose a simple cost model that treats the cost of nodes independent of other nodes.
@@
 As discussed in Section 3, our rule syntax takes one Julia expression and transforms it into another.
-This is an intuitive syntax for expressing rewrite rules, but does not fully cover all the possible rules a developer could possibly want to express.
+This is an intuitive syntax, but does not fully cover all the possible rules a developer could possibly want to express.
 For one, the rewrite rules can only be used to rewrite a single expression into another expression.
 Multi-pattern rewrite rules to match against and to rewrite multiple expressions at once are currently not supported.
-We also do not yet support rewrite rules matching against functions calls taking an unknown, variable amount of variables.
-This means that in certain cases, a developer has to rewrite multiple similar rules to match the same function with different numbers of arguments.
+We also do not yet support rewrite rules matching against function calls taking an unknown, variable amount of variables.
+A developer may hence have to rewrite multiple similar rules for the same function with different argument counts.
+Rule applicability
+Section 6 shows that certain rewrites do not always yield positive returns.
+Consequently, the burden is currently on the programmer to select the rule sets for their particular problem.
+In the future, this could be automated using an improved cost model, in combination with more detailed program analysis to gauge the effectiveness of rewrite rules.
+Alternatively, the rule applicability could be made more extensible, allowing package developers to decide programmatically when a rule can be applied.
 8. Related Work
 8.1. Equality Saturation
-Equality saturation and its use as a code optimization tool was first introduced by Tate et al.
+EqSat and its use as a code optimization tool was first introduced by Tate et al.
 (Tate et al., 2009), where it was used to optimize Java bytecode.
 Willsey et al.
-(Willsey et al., 2021) introduced algorithmic improvements for equality saturation providing asymptotic speed-ups.
-This has led to an explosion of new work exploring different applications for equality saturation, from optimizing tensor programs (Wang et al., 2020; Yang et al., 2021; Smith et al., 2021), to compiler verification (Stepp et al., 2011; Kourta et al., 2022), optimizations for specialized hardware (Ustun et al., 2022; VanHattum et al., 2021; Matsumura et al., 2023), and other purposes (Panchekha et al., 2015; Chandrakana Nandi et al., 2021).
-Metatheory.jl (Cheli, 2021b) provides a framework for equality saturation in Julia.
+(Willsey et al., 2021) introduced algorithmic improvements for EqSat providing asymptotic speed-ups.
+This has led to an explosion of new work exploring different applications for EqSat, from optimizing tensor programs (Wang et al., 2020; Yang et al., 2021; Smith et al., 2021), to compiler verification (Stepp et al., 2011; Kourta et al., 2022), optimizations for specialized hardware (Ustun et al., 2022; VanHattum et al., 2021; Matsumura et al., 2023), and other purposes (Panchekha et al., 2015; Nandi et al., 2021).
+Metatheory.jl (Cheli, 2021b) provides a framework for EqSat in Julia.
 It has already been used for code optimization on symbolic representations of Julia code to speed up PDE solvers (Gowda et al., 2022).
 That work, however, only handles pure, symbolic programs, as it operates on a symbolic representation of Julia code obtained through manual expression building or tracing through heavily restricted, straight-line code.
-In contrast to our work, that existing use of Metatheory.jl does hence not allow the presence of general control flow or side effects.
-To the best of our knowledge, by introducing type propagation in the e-graph, we are the first to apply equality saturation to a dynamic programming language.
+In contrast to our work, the existing use of Metatheory.jl does hence not allow the presence of general control flow or side effects.
+To the best of our knowledge, by introducing type propagation in the e-graph, we are the first to apply EqSat to a dynamic programming language.
 This allows rewrites to coexist and complement Julia’s multiple dispatch feature, as discussed in Section 4.
 8.2. Intermediate Representations and E-Graphs
-Cranelift, a compiler back-end for WebAssembly and Rust, is among the first to use e-graphs and equality saturation in a production-grade compiler (Fallin, 2023).
+Cranelift, a compiler back-end for WebAssembly and Rust, is among the first to use e-graphs and EqSat in a production-grade compiler (Fallin, 2023).
 While we reuse ideas and methods first introduced in the Cranelift compiler, we implemented our framework as a standalone project, not reusing any Cranelift components.
 Our code is integrated in the Julia compiler using user-land compiler extensions.
-The Cranelift compiler does not do full equality saturation, but only recognizes equalities the first time a node is inserted in the e-graph.
-This implies that it is possible for equalities to remain undiscovered, but it obviates the fixed-point loop present in full equality saturation and ensures there are no cycles in the e-graph which allows for a more efficient representation of the graph in memory and leads to easier extraction.
-We use the same e-graph representation of code as used in Cranelift, but by utilizing an extraction scheme that enforces acyclic extraction, the optimizer is able to carry out full equality saturation.
+The Cranelift compiler does not do full EqSat, but only recognizes equalities the first time a node is inserted in the e-graph.
+This implies that it is possible for equalities to remain undiscovered, but it obviates the fixed-point loop present in full EqSat and ensures there are no cycles in the e-graph which allows for a more efficient representation of the graph in memory and leads to easier extraction.
+We use the same e-graph representation of code as used in Cranelift, but by utilizing an extraction scheme that enforces acyclic extraction, the optimizer is able to carry out full EqSat.
 Whereas Cranelift and our work operate on IR that is complemented by a CFG, other work has sought to find alternative IR formats that allow representing control flow fully in the e-graph.
 Tate et al.
-(Tate et al., 2009) have introduced Program Expression Graphs (PEGs) which represent constructs such as loops as special expressions and show that PEGs can be used to do equality saturation.
+(Tate et al., 2009) have introduced Program Expression Graphs (PEGs) which represent constructs such as loops as special expressions and show that PEGs can be used to do EqSat.
+Laird et al.
+(Laird et al., 2024) raise supported LLVM IR into a functional IR that also allows control flow transformations.
 Similarly, Regionalized Value State Dependency Graphs (RVSDGs) (Bahmann et al., 2015; Reissmann et al., 2020) represent control flow as (nested) expressions.
-A research prototype already exists using RVSDGs in conjunction with equality saturation for a simple toy language (Sharp, 2022).
+A research prototype already exists using RVSDGs in conjunction with EqSat for a simple toy language (Sharp, 2022).
 Our approach remains closer to the original SSA-based IR of the source language, allowing less expensive conversion routines.
+Lastly, there are multiple works applying equality saturation on MLIR code (Merckx et al., 2025; Zayed and Dubach, 2025).
+These do not keep track of side effects and control flow in a CFG skeleton and thus can only operate on code that is pure.
+Additionally, they operate on statically typed IR in contrast to our work.
 Willsey et al.
-introduced e-class analyses to associate additional metadata with each e-class (Willsey et al., 2021).
-The metadata is kept up to date throughout the equality saturation process by potential modifications whenever new information is available, for example, when e-classes are merged.
+introduced e-class analyses, a framework to associate additional metadata with each e-class (Willsey et al., 2021).
+The metadata is kept up to date throughout the EqSat process by potential modifications whenever new information is available, for example, when e-classes are merged.
 An exemplary existing use case for an e-class analysis is constant propagation, where each e-class can optionally carry a constant value.
-To track types in the e-graph, we designed a novel e-class analysis, as discussed in Section 3.3
+Further work has shown that this technique can be combined with abstract interpretation for more complex program analyses (Coward et al., 2023b).
+To track types in the e-graph, we designed a novel e-class analysis, as discussed in Section 3.2.
 8.3. Extraction
 We based our ILP optimal extraction formulation on the work of He et al.
-(He et al., 2017).
+(He et al., 2023).
 Recent work proposes other formulations, for example by looking at e-graphs through the lens of circuits (Sun et al., 2024), or finite state automata (Y.
@@
 Other works exist that aim to improve the performance of Julia code with custom optimizations.
-Symbolics.jl (Gowda et al., 2022) can be used to transform scientific machine learning code into a symbolic representation at which point simplifications can be applied.
+Symbolics.jl (Gowda et al., 2022) can be used to transform scientific machine learning code into a symbolic representation on which simplifications can be applied.
 Finch.jl (Ahrens et al., 2024, 2023) uses a custom IR to represent iteration patterns of loop nests over sparse or structured arrays.
@@
 9. Conclusion and Future Work
-We have developed a novel system that allows Julia developers to write domain-specific rewrite rules that are automatically applied using equality saturation and that works in the presence of control flow and side effects.
+We have developed a novel system that allows Julia developers to write domain-specific rewrite rules that are automatically applied using equality saturation.
+Our system works in the presence of control flow and side effects.
 Through an e-class analysis, we keep track of the most specific type of all equivalent terms and use this information to support type-constrained rewrite rules.
@@
 We have adapted an ILP formulation for optimal, acyclic e-graph extraction to take into account value reuse from dominating statements.
-Finally, we show how this system can perform rewrites on high-level domain-specific code and how the use of equality saturation solves the phase ordering problem.
-In the future, our work can be extended to support rewriting control flow and apply rewrites that span different call depths.
+Finally, we show how this system can perform rewrites on high-level domain-specific code and how the use of equality saturation solves the phase-ordering problem.
+In the future, our work can be extended to support rewriting control flow and apply rewrites that span different call depths, by integrating function call inlining in the equality saturation procedure.
 Further improvements to our extraction scheme could also lead to more efficient code being generated.